### v1

# Contact non-squeezing in lens spaces

Abstract.
We describe some contact non-squeezing phenomena in lens spaces by defining and computing a contact capacity. This contact capacity comes from the spectral selectors constructed by Allais, Sandon and the author by the means of generating functions and Givental’s non-linear Maslov index [4]. We also discuss a potential generalization of these non-squeezing phenomena for orderable closed prequantizations.
Key words and phrases:
Contact geometry, non-squeezing, lens spaces, generating functions, Givental’s non-linear Maslov index, spectral invariants, orderability, contactomorphisms group
2020 Mathematics Subject Classification:
53D10, 57R17, 58B20
1. Introduction
1.1. Historical backgroud and main result
Gromov’s non-squeezing theorem [22] is one of the cornerstones, if not the cornestone, of symplectic geometry. Indeed, while it had been known for a long time that a symplectic manifold does not carry any local invariant, the existence of global invariants finer than the volume was not clear until Gromov’s work in [22] (see also [23, 12]). Originally stated for the standard symplectic Euclidean space, this theorem has then been generalized to all symplectic manifolds by Lalonde-McDuff [27].
For contact geometry, considered sometimes as the odd-dimensional analogue of symplectic geometry, the story is somewhat different. In this geometry, the volume is not an invariant anymore and, depending on the contact manifold considered, global invariants might not exist at all. For example, in the standard contact Euclidean space, which is the local model of contact geometry, any bounded domain can be contactly squeezed inside any arbitrary small neighborhood. Therefore, in a general contact manifold, non-squeezing phenomena cannot be expected at small scale. Nonetheless, 20 years after Gromov’s result Eliashberg, Kim and Polterovich [14] overcame this additional difficulty and proved a non-squeezing theorem at large scale in the prequantization of the standard symplectic Euclidean space. Unlike symplectic geometry, generalization of non-squeezing phenomena to other contact manifolds remain however rare. To the knowledge of the author, such phenomena have been detected only in prequantization of Liouville manifolds ([1, 6, 2, 32, 16, 8, 18]), some exotic contact spheres [35] and the unitary cotangent of the torus endowed with its canonical contact structure (see for example [13, Thm 2.4.1]).
In this paper we show that interesting non-squeezing phenomena happen also in lens spaces endowed with their standard contact structure.
Let us fix $n\in\mathbb{N}_{>0}$ a positive integer. Endow the Euclidean space $\mathbb{R}^{2n+2}$ with the coordinate functions $((x_{1},y_{1}),\cdots,(x_{n+1},y_{n+1}))$ and let $\mathbb{S}^{2n+1}:=\left\{\sum\limits_{j=1}^{n+1}x_{j}^{2}+y_{j}^{2}=1\right\}$ be the standard Euclidean sphere of radius $1$ centered at 00. Identifying $\mathbb{R}^{2n+2}$ with $\mathbb{C}^{n+1}$ via $\Phi:((x_{1},y_{1}),\cdots,(x_{n+1},y_{n+1}))\mapsto(x_{1}+iy_{1},\cdots,x_{n+1}+iy_{n+1})$ we get an action of $\mathbb{S}^{1}:=\{e^{i\theta}\ |\ \theta\in\mathbb{R}\}\subset\mathbb{C}$ on $\mathbb{S}^{2n+1}\subset\mathbb{C}^{n+1}$ given by multiplication. For all integer $k\geq 1$, the lens space $L_{k}^{2n+1}$ is the quotient space $\mathbb{S}^{2n+1}/\mathbb{Z}_{k}$ where $\mathbb{Z}_{k}:=\mathbb{Z}/k\mathbb{Z}\simeq\{e^{\frac{2i\pi m}{k}}\ |\ m\in[0,k-1]\cap\mathbb{N}\}$ is identified with the $k$th roots of unity. We denote by $\pi_{1}^{k}:\mathbb{S}^{2n+1}\to L_{k}^{2n+1}$ the covering map. It carries a natural contact form $\alpha_{k}$ defined as follow: $(\pi^{k}_{1})^{*}\alpha_{k}=\alpha_{1}$ where $\alpha_{1}=i_{\mathbb{S}^{2n+1}}^{*}\left(\sum\limits_{j=1}^{n+1}(x_{j}dy_{j}-y_{j}dx_{j})\right)$ and $i_{\mathbb{S}^{2n+1}}:\mathbb{S}^{2n+1}\hookrightarrow\mathbb{R}^{2n+2}$ is the inclusion.
Finally we identify the complex projective space $\mathbb{C}\mathrm{P}^{n}$ with the quotient space $\mathbb{S}^{2n+1}/\mathbb{S}^{1}$ and denote by $\pi_{1}:\mathbb{S}^{2n+1}\to\mathbb{C}\mathrm{P}^{n}$ the projection. We consider on $\mathbb{C}\mathrm{P}^{n}$ the only symplectic form $\omega$ that satisfies $\pi_{1}^{*}\omega:=\mathrm{d}\alpha_{1}$. Note that $\omega=2\omega_{\mathrm{FS}}$, where $\omega_{\mathrm{FS}}$ denotes the standard Fubini-Study form, and that $\pi_{k}^{*}\omega=\mathrm{d}\alpha_{k}$, where $\pi_{k}:L_{k}^{2n+1}\to\mathbb{C}\mathrm{P}^{n}$ is the natural projection.
$M$ being a smooth manifold and $\xi$ a contact distribution on $M$, we say that a subset $U_{1}\subset M$ cannot be contactly squeezed inside another subset $U_{2}\subset M$ if for all contactomorphisms $\phi\in\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ isotopic to the identity through a compactly supported contact isotopy, $\phi(U_{1})$ does not lie inside $U_{2}$, i.e. $\phi(U_{1})\cap U_{2}\neq\phi(U_{1})$.
Denoting by $\mathbb{B}^{2n}(r):=\{((x_{1},y_{1}),\cdots,(x_{n},y_{n}))\in\mathbb{R}^{2n}\ |\ \sum\limits_{j=1}^{n}x_{j}^{2}+y_{j}^{2}<r^{2}\}$ the standard open Euclidean ball of radius $r>0$ in $\mathbb{R}^{2n}$ centered at 00 and by $\omega_{0}:=\sum\limits_{j=1}^{n}dx_{j}\wedge dy_{j}$ the standard symplectic form of $\mathbb{R}^{2n}$ we will prove the following:
Theorem 1.1.
Let $\Psi_{1}:(\mathbb{B}^{2n}(R_{1}),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ and $\Psi_{2}:(\mathbb{B}^{2n}(\lambda R_{2}),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ be two symplectic embeddings for some $R_{1},\ R_{2}>0$, $\lambda>\sqrt{3}$ and let $a_{1}\in(0,R_{1}],\ a_{2}\in(0,R_{2}]$ be two positive real numbers. For any integer $k\geq 2$, if there exists $j\in\mathbb{N}_{>0}$ such that
$\pi a_{1}^{2}>\frac{2\pi}{k}\cdot j\geq\pi a_{2}^{2},$
then $B^{1}_{k}(a_{1})$ cannot be contactly squeezed inside $B^{2}_{k}(a_{2})$, where by definition $B^{i}_{k}(a_{i}):=\pi_{k}^{-1}(\Psi_{i}(\mathbb{B}^{2n}(a_{i})))\subset(L_{k}^{2n+1},\xi_{k})$ for all $i\in\{1,2\}$.
Remark 1.2.
  1. (1)
    It might be that Theorem 1.1 remains true for some $\lambda\leq\sqrt{3}$. However, for the purpose of our proof, $\lambda$ has to be strictly greater than $1$ since, as we will see in Subsection 5.2, we need to displace $\mathbb{B}^{2n}(a_{2})$ from itself inside $\mathbb{B}^{2n}(\lambda R_{2})$.
  2. (2)
    These non-squeezing phenomena do not happen if we forget about the contact structure. Indeed it is easy to see that if $R_{1}=\lambda R_{2}=:R$ and $\Psi_{1}=\Psi_{2}=:\Psi$ then for all $r_{1},\ r_{2}\in(0,R]$, there exists a diffeomorphism of $L_{k}^{2n+1}$ which is isotopic to the the identity - through diffeomorphisms - and that sends $B_{k}(r_{1})$ inside $B_{k}(r_{2})$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$ for all $r\in(0,R]$.
  3. (3)
    In [26, Thm 1 and Prop 2.11] [^1] The conventions used in [26] are different: they consider the symplectic form $\widetilde{\omega}:=\frac{1}{2\pi}{\omega}$ and show that there exists an embedding of $(\widetilde{\mathbb{B}^{2n}}(1),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\widetilde{\omega})$, where $\widetilde{\mathbb{B}^{2n}}(a):=\mathbb{B}^{2n}(\sqrt{\frac{a}{\pi}})$ Karshon and Tolman construct a symplectic embedding $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$, for $R=\sqrt{2}$, and show that it is optimal in the sense that such symplectic embeddings do not exist for $R>\sqrt{2}$ (see also [28, Thm 1]). Considering this optimal embedding, one can see that non-squeezing phenomena appear for all lens spaces, even for the real projective space $(\mathbb{R}\mathrm{P}^{2n+1},\xi):=(L_{2}^{2n+1},\xi_{2})$, since $\pi\left(\sqrt{2}\right)^{2}>\frac{2\pi}{2}$.
We will prove Theorem 1.1 by constructing a contact capacity defined on open subsets of $(L_{k}^{2n+1},\xi_{k})$ and by estimating this contact capacity for the lift of Darboux balls. This contact capacity comes from the spectral selectors defined on $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ in [4].
1.2. Discussion
It is interesting to note that the $\alpha_{k}$-Reeb flow in $L_{k}^{2n+1}$ is periodic with period $\frac{2\pi}{k}$ for all integers $k\geq 1$. Indeed, the periodicity of the Reeb flow often plays a crucial role in the existence of non trivial global invariants and this article is no exception to this fact (see for example Remark 2.4 and Lemma 6.2). Furthermore, the global invariants one can deduce from Theorem 1.1 naturally take value in the discrete set consisting in multiples of this period. Similarly, in the contact manifold $\left(\mathbb{R}^{2n}\times\mathbb{R}/\mathbb{Z},\alpha_{0}:=\mathrm{d}z-\frac{1}{2}\sum\limits_{i=1}^{n}(x_{i}dy_{i}-y_{i}dx_{i})\right)$ where the first contact non-squeezing phenomena was described by Eliashberg, Kim and Polterovich in [14], the $\alpha_{0}$-Reeb flow is $1$-periodic, and they show that it is impossible to contactly squeeze $\mathbb{B}^{2n}(A)\times\mathbb{R}/\mathbb{Z}$ inside $\mathbb{B}^{2n}(a)\times\mathbb{R}/\mathbb{Z}$ if there exists an integer $j\in\mathbb{N}_{>0}$ so that $A>j\geq a$.
The same kind of phenomenon can be observed when studying bi-invariant metrics on the group $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ or its universal cover $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(M,\xi)$. While Sandon in [31] constructs an unbounded bi-invariant metric on $\mathrm{Cont}_{0}^{\mathrm{c}}(\mathbb{R}^{2n}\times\mathbb{R}/\mathbb{Z},\ker\alpha_{0})$ taking integer values, the authors in [4] construct an unbounded bi-invariant metric on $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ taking values in $\frac{2\pi}{k}\mathbb{N}$ [^2] It is shown in [17] that, for any contact manifold $(M,\xi)$, a bi-invariant metric on $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ (resp. $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(M,\xi)$) has to induce the discrete topology on $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ (resp. on $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(M,\xi)\setminus\pi_{1}(\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi))$).
What we have just described should be nuanced: beeing a contact manifold admitting a periodic Reeb flow is not a sufficient condition to have rigidity results and existence of global invariants. In fact, the $\alpha_{1}$-Reeb flow of the standard contact sphere $(\mathbb{S}^{2n+1},\ker\alpha_{1})$ is periodic but any two non-empty open domains that are not the whole sphere can be contactly squeezed inside each other. In the same way, it is shown in [17] that, for $n\in\mathbb{N}_{>0}$, the groups $\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\ker\alpha_{1})$ and $\widetilde{\mathrm{Cont}_{0}}(\mathbb{S}^{2n+1},\ker\alpha_{1})$ are bounded, i.e. any bi-invariant metric on these groups has to be bounded.
It seems that another important property that a contact manifold should have so that one could expect meaningful invariant measurements is to be orderable (see Section 6 for a definition). For $n\in\mathbb{N}_{>0}$ and $k\geq 2$, the spheres $(\mathbb{S}^{2n+1},\ker\alpha_{1})$ and the lens spaces $(L_{k}^{2n+1},\ker\alpha_{k})$ are perfect representative cases since the sphere is not orderable [14] while the lens space is [21], and, as we have just mentionned, interesting contact measurements exist for the latter but not for the former. In the same direction, a more general observation can be made for bi-invariant metrics on contactomorphisms groups. On the one hand, to the knowledge of the author, the only examples in the litterature of closed contact manifolds $(M,\xi)$ admitting unbounded bi-invariant metrics on $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ are the contact manifolds $(M,\xi)$ that are orderable and that admit a periodic Reeb flow [17, 10, 21, 4]. On the other hand, for a large family of closed non-orderable contact manifolds $(M,\xi)$ Courte and Massot show that $\mathrm{Cont}_{0}(M,\xi)$ and $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ have to be bounded [11].
That being so, we discuss in Section 6 a potential generalization of Theorem 1.1 to orderable prequantizations of closed symplectic manifolds which are particular cases of orderable closed contact manifold with periodic Reeb flow.
Actually, while convenient but not sufficient, the existence of a periodic Reeb flow is not even a necessary condition to have contact global invariants. For example, the unitary cotangent of the torus $\mathbb{P}^{+}T^{*}\mathbb{T}^{n}$ endowed with its canonical contact structure $\xi_{\mathrm{can}}$ does not admit any periodic Reeb flow when $n>1$ (see for example [7]). But, as briefly mentionned above, one can use an interesting contact global invariant for domains in $(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$, which is called the shape333unlike the other invariant measurements discussed so far, the shape can be used to construct “non-discrete” measurements, to deduce non-squeezing phenomena and other rigidity results [13, 6, 29, 15]. It would therefore be interesting to investigate whether the groups $\widetilde{\mathrm{Cont}_{0}}(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$ and $\ \mathrm{Cont}_{0}(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$ are bounded or not (see also [3, Prop. 4.14, Rem. 4.15]).
Organization of the paper
In Section 2 we recall briefly the construction and main properties of the spectral selectors constructed in [4]. In Section 3 we define the above mentionned contact capacity for open domains of $(L_{k}^{2n+1},\xi_{k})$ coming from the spectral selectors. In this same Section 3 we state - without prooving - some properties of this contact capacity (see Theorem 3.1) which allow us to give the proof of Theorem 1.1. In Section 4 we show a energy-capacity inequality satisfied by the contact capacity. In Section 5 we prove Theorem 3.1 by estimating and computing the contact capacity of domains that are lifts of Darboux balls using the energy-capacity inequality of Section 4 and some explicit constructions. Finally, we discuss in Section 6 on a possible generalization of this result to orderable closed prequantizations.
Acknowledgment
The author thanks Alberto Abbondandolo, Johanna Bimmermann, Stefan Nemirovski and Sheila Sandon for interesting and stimulating discussions.
The author was partially supported by the Deutsche Forschungsgemeinschaft under the Collaborative Research Center SFB/TRR 191 - 281071066 (Symplectic Structures in Geometry, Algebra and Dynamics).
2. Spectral Selectors
2.1. Conventions
Let $(M,\xi)$ be a closed coorientable contact manifold. We say that a $[0,1]$-family of contactomorphisms $\{\phi_{t}\}_{t\in[0,1]}\subset\mathrm{Cont}(M,\xi)$, that we abbreviate with $(\phi_{t})$, is a contact isotopy if the map $[0,1]\times M\to M$, $x\mapsto\phi_{t}(x)$ is smooth. We denote by $\mathrm{Cont}_{0}(M,\xi)$ the group of contactomorphisms $\phi$ that are contact isotopic to the identity: there exists a contact isotopy $(\phi_{t})$ such that $\phi_{0}=\mathrm{id}$ and $\phi_{1}=\phi$. We endow $\mathrm{Cont}_{0}(M,\xi)$ with the $C^{1}$-topology, denote by $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ its universal cover and $\Pi:\widetilde{\mathrm{Cont}_{0}}(M,\xi)\to\mathrm{Cont}_{0}(M,\xi)$ the projection. We often see $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ as equivalence classes of contact isotopies starting at the identity where two isotopies represent the same equivalence class if they are homotopic relatively to endpoints.
A point $x\in M$ is said to be a discriminant of a contactomorphism $\phi\in\mathrm{Cont}_{0}(M,\xi)$ if $\phi(x)=x$ and $(\phi^{*}\alpha)_{x}=\alpha_{x}$, for some - and thus all - contact form $\alpha$ supporting $\xi$, i.e. $\ker\alpha=\xi$.
Let $\alpha$ be a contact form supporting $\xi$, i.e. $\ker\alpha=\xi$. We say that a point $x\in M$ is an $\alpha$-translated point of $\phi\in\mathrm{Cont}_{0}(M,\xi)$ if there exists $T\in\mathbb{R}$ so that $x$ is a discriminant point of $\phi_{\alpha}^{-T}\circ\phi$, where $\phi_{\alpha}^{s}$ denotes the Reeb flow at time $s\in\mathbb{R}$ of $\alpha$, i.e. the flow of the vector field $R_{\alpha}$ uniquely defined by the two equations $\alpha(R_{\alpha})\equiv 1$ and $\mathrm{d}\alpha(R_{\alpha},\cdot)\equiv 0$. For any $\phi\in\mathrm{Cont}_{0}(M,\xi)$ we denote and define its $\alpha$-spectrum by
$\mathrm{Spec}^{\alpha}(\phi):=\left\{t\in\mathbb{R}\ |\ \phi_{\alpha}^{-t}\circ\phi\text{ has a discriminant point}\right\}.$
Similarly, the $\alpha$-spectrum $\mathrm{Spec}^{\alpha}(\widetilde{\varphi})$ of any $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ is by definition
$\mathrm{Spec}^{\alpha}(\widetilde{\varphi}):=\mathrm{Spec}^{\alpha}\left(\Pi(\widetilde{\varphi})\right).$
Definition 2.1.
We say that a map $c:\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)\to\mathbb{R}$ is a $\alpha$-spectral selector if it is $C^{1}$-continuous and if $c(\widetilde{\varphi})\in\mathrm{Spec}^{\alpha}(\widetilde{\varphi})$.
In [4] the authors construct a $\alpha_{k}$-spectral selector $c_{k}:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\ker\alpha_{k})\to\mathbb{R}$ for any $n\in\mathbb{N}_{>0}$ and $k\geq 2$ (denoted in [4] by $c_{0}$ or $c_{+}$). Its construction, that we present briefly below, is based on generating functions techniques and Givental’s non-linear Maslov index [21, 20]. From now on we fix $n\in\mathbb{N}_{>0}$ and an integer $k\geq 2$.
2.2. Givental’s non-linear Maslov index
The Givental’s non-linear Maslov index $\mu:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{Z}$ is a quasimorphism that detects discriminant points in the following sense: if $\{\widetilde{\varphi}_{t}\}_{t\in[0,1]}\subset\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ is a smooth path, i.e. $(\phi_{t}:=\Pi(\widetilde{\varphi}_{t}))$ is a contact isotopy, such that $\mu(\widetilde{\varphi}_{0})\neq\mu(\widetilde{\varphi}_{1})$, then there exists $t_{0}\in[0,1]$ such that $\Pi(\widetilde{\varphi}_{t_{0}})$ has a discriminant point. Denoting by $\widetilde{\phi_{\alpha_{k}}^{T}}:=[(\phi_{\alpha_{k}}^{tT})]\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$, for $T\in\mathbb{R}$, it is shown moreover in [20, 21, 4] that $\mu$ has the following properties.
Lemma 2.2 ([20, 21, 4]).
For all $\widetilde{\varphi},\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
  1. (1)
    the map $\mathbb{R}\to\mathbb{Z}$, $T\mapsto\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})$ is lower semi-continuous
  2. (2)
    if $\mu(\widetilde{\varphi})$ is even then $\mu(\widetilde{\psi}\cdot\widetilde{\varphi})\leq\mu(\widetilde{\psi})+\mu(\widetilde{\varphi})$
  3. (3)
    $\mu(\widetilde{\phi_{\alpha_{k}}^{T}})=2(n+1)\left\lceil\frac{T}{2\pi}\right\rceil$ for all $T\in\mathbb{R}$
  4. (4)
    $\mu(\widetilde{\varphi})\leq\mu(\widetilde{\psi})$ if $\widetilde{\varphi}\preceq\widetilde{\psi}$,
where $\preceq$ is the bi-invariant binary relation introduced by Eliashberg-Polterovich in [15], i.e. $\widetilde{\varphi}\preceq\widetilde{\psi}$ if $\widetilde{\psi}\cdot\widetilde{\varphi}^{-1}$ can be represented by a contact isotopy $(\phi_{t})$ such that $\alpha_{k}\left(\frac{d}{dt}\phi_{t}\right)\geq 0$.
Let us give a brief overview of its construction. To do consider $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ and $(\phi_{t})$ any contact isotopy representing it.
Firstly, let us denote by $(\overline{\phi_{t}})\subset\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\xi_{1})$ the lifted $\mathbb{Z}_{k}$-equivariant contact isotopy starting at the identity associated to the contact isotopy $(\phi_{t})$, i.e. $\pi_{1}^{k}\circ\overline{\phi_{t}}=\phi_{t}\circ\pi_{1}^{k}$ for all $t\in[0,1]$ and $\overline{\phi_{0}}=\mathrm{id}$. We want to construct a path of $C^{1}$-functions $(f_{t}:L_{k}^{2(n+N)+1}\to\mathbb{R})$, for some integer $N\geq 0$, so that for all $t\in[0,1]$
$\{x\in L_{k}^{2(n+N)+1}\ |\ \mathrm{d}_{x}f_{t}=0,\ f_{t}(x)=0\}\overset{1-1}{\leftrightarrow}\{\pi_{1}^{k}(x)\in L_{k}^{2n+1}\ |\ x\text{ is a disc. point of }\overline{\phi}_{t}\},$
i.e. there is a one-to-one correspondance between the critical points of $f_{t}$ with critical value 0 and the discriminant points of $\phi_{t}$ that lift to a $\mathbb{Z}_{k}$-family of discriminant points of $\overline{\phi}_{t}$. This path of function $(f_{t})$ is constructed in the following way. Identifying the symplectization of $(\mathbb{S}^{2n+1},\xi_{1})$ with $(\mathbb{R}^{2n+2}\setminus\{0\},\omega_{0})$, the $\mathbb{Z}_{k}$-equivariant contact isotopy $(\overline{\phi}_{t})\subset\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\xi_{1})$, can itself be uniquely lifted to a $\mathbb{Z}_{k}\times\mathbb{R}_{>0}$-equivariant Hamiltonian isotopy $(\psi_{t})\subset\mathrm{Ham}(\mathbb{R}^{2n+2}\setminus\{0\},\omega_{0})$
$\psi_{t}(z)=e^{-\frac{1}{2}g_{t}(z/\|z\|)}\|z\|\overline{\phi}_{t}\left(\frac{z}{\|z\|}\right),\quad\text{ for all }t\in[0,1],\ z\in\mathbb{R}^{2n+2}\setminus\{0\},$
where $g_{t}:\mathbb{S}^{2n+1}\to\mathbb{R}$ is the $\alpha_{1}$-conformal factor of $\overline{\phi}_{t}$, i.e. $\overline{\phi}_{t}^{*}\alpha_{1}=e^{g_{t}}\alpha_{1}$. Extending $\psi_{t}$ to the whole $\mathbb{R}^{2n+2}$ by putting $\psi_{t}(0)=0$ the authors in [21, Prop 2.2 and Prop 2.14] show the existence of an integer $m\in\mathbb{N}$ and a path of $C^{1}$-functions $(F_{t}:\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}\to\mathbb{R})$, where $2N=(2n+2)\cdot 2m$, such that denoting by $\Sigma_{F_{t}}:=\{(x,v)\in\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}\ |\ \partial_{v}F(x,v)\equiv 0\in T^{*}\mathbb{R}^{2n}\}$ the fiber critical points of $F_{t}$ we have444 in the following $\mathbb{R}^{2j}$ is identified with $\mathbb{C}^{j}$, for all $j\in\mathbb{N}_{>0}$, as in the introduction and $T^{*}\mathbb{R}^{2n+2}$ with $T\mathbb{R}^{2n+2}\simeq\mathbb{R}^{2n+2}\times\mathbb{R}^{2n+2}$ using the standard Euclidean metric
  1. (1)
    $i_{F_{t}}:\Sigma_{F_{t}}\to T^{*}\mathbb{R}^{2n+2},\ (x,v)\hookrightarrow\left(x,\partial_{x}F_{t}(x,v)\right)$ is injective and moreover
    $i_{F_{t}}(\Sigma_{F_{t}})=\tau(\mathrm{graph}(\psi_{t}))$
    where $\text{graph}(\psi_{t}):=\{(x,\psi_{t}(x))\ |\ x\in\mathbb{R}^{2n+2}\}$ and $\tau(z,Z)=\left(\frac{z+Z}{2},i(z-Z)\right)$ for all $z,Z\in\mathbb{R}^{2n+2}\times\mathbb{R}^{2n+2}.$
  2. (2)
    $F_{t}(\lambda e^{\frac{2i\pi j}{k}}z)=\lambda^{2}F_{t}(z)$ for all $\lambda>0$, $z\in\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}$ and $j\in\mathbb{N}$
  3. (3)
    $F_{0}:\mathbb{R}^{2n+2}\times(\mathbb{R}^{2n+2})^{2m}\to\mathbb{R}$, $(z,(Z_{1},\cdots,Z_{2m}))\mapsto-2\sum\limits_{j=1}^{m}\left<Z_{2j}-z,i(Z_{2j-1}-z)\right>$.
The path of functions $(F_{t})$ induces the desired path $(f_{t}:L_{k}^{2(n+N)+1}\to\mathbb{R})$ and $(F_{t})$ is called a based family of generating functions associated to $(\phi_{t})$.
Secondly, suppose that $k=p\in\mathbb{N}_{>0}$ is a prime number. If $(\phi^{\prime}_{t})$ is another contact isotopy in $\mathrm{Cont}_{0}(L_{p}^{2n+1},\xi_{p})$ representing the same element $\widetilde{\varphi}$ and if $(f_{t}:L_{p}^{2N+1}\to\mathbb{R})$ and $(f_{t}^{\prime}:L_{p}^{2N^{\prime}+1}\to\mathbb{R})$ are two paths of functions that are induced by two based families of generating functions associated respectively to $(\phi_{t})$ and $(\phi^{\prime}_{t})$, then it is proven in [21, Prop. 2.20 and Section 4] that
$\operatorname{ind}(\{f_{0}\leq 0\})-\operatorname{ind}(\{f_{1}\leq 0\})=\operatorname{ind}(\{f^{\prime}_{0}\leq 0\})-\operatorname{ind}(\{f_{1}^{\prime}\leq 0\}).$
Here the index $\operatorname{ind}(A)$ of an open subset $A\subset L_{p}^{j}$, for any integer $j\geq 3$, is the dimension over the field $\mathbb{Z}_{p}$ of the image of $i_{A}^{*}:H^{*}(L_{p}^{j},\mathbb{Z}_{p})\to H^{*}(A,\mathbb{Z}_{p})$, where $i_{A}$ denotes the inclusion and $i_{A}^{*}$ the induced map on singular homology, and the index of a closed subset $F$ is by definition $\operatorname{ind}(F):=\min\{\operatorname{ind}(A)\ |\ A\text{ open and }F\subset A\}$. The Givental’s non-linear Maslov index of $\widetilde{\varphi}$ is then defined to be
$\mu(\widetilde{\varphi}):=\operatorname{ind}(\{f_{0}\leq 0\})-\operatorname{ind}(\{f_{1}\leq 0\}).$
Finally, for a general integer $k\geq 2$, denoting by $p\in\mathbb{N}_{>0}$ the smallest prime dividing $k$, we define $\mu(\widetilde{\varphi})$ by
$\mu(\widetilde{\varphi}):=\mu(\mathcal{L}_{k}^{p}(\widetilde{\varphi})),$
where $\mathcal{L}_{k}^{p}:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\widetilde{\mathrm{Cont}_{0}}(L_{p}^{2n+1},\xi_{p})$ denotes the natural lift map.
2.3. The spectral selector on lens spaces
The $\alpha_{k}$-spectral introduced in [4] is defined by
$c_{k}:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{R}\quad\quad\widetilde{\varphi}\mapsto\inf\left\{T\in\mathbb{R}\ |\ \mu\left(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi}\right)\leq 0\right\}.$
Before stating its main properties we introduce some notations. For any $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ we consider the following set
$\overline{\mathrm{Spec}^{\alpha_{k}}}(\widetilde{\varphi}):=\mathrm{Spec}^{\alpha_{1}}(\overline{\widetilde{\varphi}})\subset\mathrm{Spec}^{\alpha_{k}}(\widetilde{\varphi})$
where $\overline{\widetilde{\varphi}}\in\widetilde{\mathrm{Cont}_{0}}(\mathbb{S}^{2n+1},\xi_{1})$ denotes the $\mathbb{Z}_{k}$-equivariant lift of $\widetilde{\varphi}$. For any positive number $T>0$ and any $x\in\mathbb{R}$ we denote by $\lceil x\rceil_{T}:=T\left\lceil\frac{x}{T}\right\rceil$ the smallest multiple of $T$ greater or equal to $x$.
Proposition 2.3 ([4]).
The map $c_{k}:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{R}$ is a $\alpha_{k}$-spectral selector that moreover satisfies for all $\widetilde{\varphi},\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
  1. (1)
    $c_{k}(\mathrm{id})=0$
  2. (2)
    $\widetilde{\varphi}\preceq\widetilde{\psi}$ implies that $c_{k}(\widetilde{\varphi})\leq c_{k}(\widetilde{\psi})$
  3. (3)
    $\mathrm{id}\preceq\widetilde{\varphi}$ and $\widetilde{\varphi}\neq\mathrm{id}$ implies that $c_{k}(\widetilde{\varphi})>0$
  4. (4)
    $c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi})\leq c_{k}(\widetilde{\psi})+\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$
  5. (5)
    for any contact isotopy $(\phi_{t})$ representing $\widetilde{\varphi}$
    $\int_{0}^{1}\min\alpha_{k}\left(\frac{d}{dt}\phi_{t}\right)\ \mathrm{d}t\leq c_{k}(\widetilde{\varphi})\leq\int_{0}^{1}\max\alpha_{k}\left(\frac{d}{dt}\phi_{t}\right)\ \mathrm{d}t$
  6. (6)
    $c_{k}(\widetilde{\varphi})\in\overline{\mathrm{Spec}^{\alpha_{k}}}(\widetilde{\varphi})$
  7. (7)
    $\lceil c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}^{-1})\rceil_{\frac{2\pi}{k}}=\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$.
Remark 2.4.
Lemma 6.2 below states that the conjugation invariance property of the map $\widetilde{\varphi}\mapsto\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$ is actually a consequence of the $\frac{2\pi}{k}$-periodicity of the $\alpha_{k}$-Reeb flow and the $\alpha_{k}$-spectrality of $c_{k}$. This was already noticed by Sandon [32] in the case of the prequantization of the standard symplectic Euclidean space.
Before concluding this section it is worth noticing that we have a stronger triangular inequality and invariance by conjugation for elements in
[…]
Proposition 2.5.
For any $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{S}^{1}}(L_{k}^{2n+1},\xi_{k})$, $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
  1. (1)
    $c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi})\leq c_{k}(\widetilde{\psi})+c_{k}(\widetilde{\varphi})$
  2. (2)
    $c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}^{-1})=c_{k}(\widetilde{\varphi})$.
Remark 2.6.
For the purpose of this paper the second statement of Proposition 2.5 is not relevant.
To prove Proposition 2.5, let us recall that $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is said to be $\alpha_{k}$-nondegenerate if for all $\alpha_{k}$-translated points $x\in L_{k}^{2n+1}$ of $\phi$ there does not exist any $X\in T_{x}L_{k}^{2n+1}\setminus\{0\}$ and $T\in\mathbb{R}$ such that
$\mathrm{d}_{x}(\phi_{\alpha_{k}}^{-T}\circ\phi)(X)=X\text{ and }\mathrm{d}_{x}g(X)=0$
where $g:L_{k}^{2n+1}\to\mathbb{R}$ is the $\alpha_{k}$-conformal factor of $\phi$, i.e. $\phi^{*}\alpha_{k}=e^{g}\alpha_{k}$. The next Lemma is also contained in [4, Thm 1.1].
Lemma 2.7.
Let $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$. If $\Pi(\widetilde{\varphi})$ is $\alpha_{k}$-nondegenerate then $\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})=0$ where $T:=c_{k}(\widetilde{\varphi})$.
Proof of Proposition.
2.5
  1. (1)
    Let us first suppose that $\Pi(\widetilde{\varphi})$ is $\alpha_{k}$-nondegenerate and denote by $T:=c_{k}(\widetilde{\varphi})$, $T^{\prime}:=c_{k}(\widetilde{\psi})$. Then by Lemma 2.7 we have that $\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})=0$ is even. Therefore, since $\widetilde{\psi}$ commutes with $\widetilde{\phi_{\alpha_{k}}^{s}}$ for all $s\in\mathbb{R}$ we have
    $\mu(\widetilde{\phi_{\alpha_{k}}^{-(T^{\prime}+T)}}\cdot\widetilde{\psi}\cdot\widetilde{\varphi})=\mu(\widetilde{\phi_{\alpha_{k}}^{-T^{\prime}}}\cdot\widetilde{\psi}\cdot\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})$ $\leq\mu(\widetilde{\phi_{\alpha_{k}}^{-T^{\prime}}}\cdot\widetilde{\psi})+\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})$
    $\leq\mu(\widetilde{\phi_{\alpha_{k}}^{-T^{\prime}}}\cdot\widetilde{\varphi})\leq 0,$
    where the first inequality follows from the 2nd point of Lemma 2.2 and the last inequality from the 1st property of Lemma 2.2. This implies that
    $c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi})\leq c_{k}(\widetilde{\psi})+c_{k}(\widetilde{\varphi})$
    which finishes the proof in the case where $\Pi(\widetilde{\varphi})$ is non degenerate.
    The general case can now be deduced by the $C^{1}$-continuity of $c_{k}$ and the $C^{1}$-density, inside $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$, of the set of elements $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ such that $\Pi(\widetilde{\psi})$ is $\alpha_{k}$-nondegenerate (it is a direct consequence of [4, Lemma 3.1] and the fact that $\Pi:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is a local homeomorphism when endowing both spaces with the $C^{1}$-topology).
  2. (2)
    Let $(\psi_{t})$ be a contact isotopy representing $\widetilde{\psi}$ such that $\phi_{\alpha_{k}}^{s}\circ\psi_{t}=\psi_{t}\circ\phi_{\alpha_{k}}^{s}$ for all $t\in[0,1]$ and $s\in\mathbb{R}$. This implies that $\psi_{T}^{*}\alpha_{k}=\alpha_{k}$, for all $T\in[0,1]$, and that
    $\mathrm{Spec}^{\alpha_{k}}(\widetilde{\psi}_{T}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{T}^{-1})=\mathrm{Spec}^{\alpha_{k}}(\widetilde{\varphi})$
    where $\widetilde{\psi}_{T}:=[(\psi_{tT})]$. Moreover it is proved in [4, Lemma 3.2] that $\mathrm{Spec}^{\alpha_{k}}(\widetilde{\varphi})$ is nowhere dense which implies that the map $[0,1]\to\mathbb{R},\ T\mapsto c_{k}(\widetilde{\psi}_{T}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{T}^{-1})\in\mathrm{Spec}^{\alpha_{k}}(\widetilde{\varphi})$ is constant, because continuous, and thus
    $c_{k}(\widetilde{\varphi})=c_{k}(\widetilde{\psi}_{0}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{0}^{-1})=c_{k}(\widetilde{\psi}_{1}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{1}^{-1})=c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}^{-1}).$
Remark 2.8.
A $\mathbb{S}^{1}$-equivariant contactomorphism $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is always $\alpha_{k}$-degenerate since its $\alpha_{k}$-conformal factor is constantly equal to 00 and $(\phi_{\alpha_{k}}^{-T}\circ\phi)^{*}R_{\alpha_{k}}=R_{\alpha_{k}}$ for all $T\in\mathbb{R}$.
3. Contact capacity
For any non-empty open set $U\subset L_{k}^{2n+1}$ we denote by $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})$ the subgroup of $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ whose elements can be represented by contact isotopies $(\phi_{t})$ supported in $U$, i.e. $\overline{\underset{t\in[0,1]}{\bigcup}\mathrm{Supp}(\phi_{t})}\subset U$ where $\mathrm{Supp}(\phi_{t}):=\overline{\{x\in M\ |\ \phi_{t}(x)\neq x\}}$. Let $C_{k}:\{\text{non-empty open sets of }L_{k}^{2n+1}\}\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ be the map:
$U\mapsto C_{k}(U):=\sup\{c_{k}(\widetilde{\varphi})\ |\ \widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})\}.$
Theorem 3.1.
For all non-empty open sets $U,V\subset L_{k}^{2n+1}$ and $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$
  1. (1)
    (monotonicity) $C_{k}(U)\leq C_{k}(V)$ if $U\subset V$
  2. (2)
    (invariance by conjugation) $\lceil C_{k}(U)\rceil_{\frac{2\pi}{k}}=\lceil C_{k}(\phi(U))\rceil_{\frac{2\pi}{k}}$
  1. (3)
    (lower bound) if $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ is a symplectic embedding for some $R>0$ then $C_{k}(B_{k}(r))\geq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$,
  2. (4)
    (upper bound) if $\Psi:(\mathbb{B}^{2n}(\lambda R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ is a symplectic embedding for some $R>0$ and $\lambda>\sqrt{3}$ then $C_{k}(B_{k}(r))\leq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
The monotonicity property of $C_{k}$ in Theorem 3.1 follows by definition. The invariance by conjugation of $C_{k}$ follows from the invariance by conjugation of the map $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{R},\ \widetilde{\varphi}\mapsto\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$ (see Proposition 2.3). To give the complete proof of Theorem 3.1 it thus remains to prove the last two estimates. Let us admit them for the moment and deduce the proof of Theorem 1.1:
Proof of Theorem 1.1.
Suppose by contradiction that there exists $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ such that $\phi(B^{1}_{k}(a_{1}))\subset B^{2}_{k}(a_{2})$. Then, using the 1st and 4th point of Theorem 3.1, we have on one side
$C_{k}(\phi(B^{1}_{k}(a_{1})))\leq C_{k}(B^{2}_{k}(a_{2}))\leq\pi a_{2}^{2}\leq\frac{2\pi}{k}j.$ (1)
But on the other side using the 2nd and the 3rd point we get that
$\lceil C_{k}(\phi(B^{1}_{k}(a_{1})))\rceil_{\frac{2\pi}{k}}=\lceil C_{k}(B^{1}_{k}(a_{1}))\rceil_{\frac{2\pi}{k}}\geq\left\lceil\pi a_{1}^{2}\right\rceil_{\frac{2\pi}{k}}>\frac{2\pi}{k}j$
which contradicts (1) and finishes the proof of Theorem 1.1. ∎
We give the proof of the last two estimates of Theorem 3.1 in Section 5. But before, in the next Section 4 we prove a energy-capacity inequality satisfied by $C_{k}$ which will be crucial to prove the lower bound stated in Theorem 3.1.
4. energy-capacity Inequality
Let $(M,\ker\alpha)$ be a closed contact manifold. For any subset $U\subset M$ we denote by $\mathcal{O}_{\alpha}(U):=\underset{t\in\mathbb{R}}{\bigcup}\phi_{\alpha}^{t}(U)$ its $\alpha$-orbit.
Definition 4.1.
A diffeomorphism $\phi$ of $M$ $\alpha$-displaces a subset $U\subset M$ if $\phi$ displaces the $\alpha$-orbit of $U$, i.e. $\phi(\mathcal{O}_{\alpha}(U))\bigcap\mathcal{O}_{\alpha}(U)=\emptyset.$
Remark 4.2.
  1. (1)
    If $(M,\xi)$ is a closed cooriented contact manifold, the Reeb flow of a generic contact form supporting $\xi$ has a dense orbit [9, Theorem C.4]. Hence, for such contact form $\alpha$, a diffeomorphism $\phi$ cannot displace the closure of the $\alpha$-orbit of any non-empty open set $U$, i.e. $\phi(\overline{\mathcal{O}_{\alpha}(U)})\cap\overline{\mathcal{O}_{\alpha}(U)}\neq\emptyset$.
  2. (2)
[…]
  3. (3)
    The first two remarks has to be put in contrast with the following fact: it has been shown in [33] that any contact manifold $(M,\xi)$ admits a contact form $\alpha$ and a non-empty open set $U\subset M$ where the $\alpha$-Reeb flow is $1$-periodic on $U$.
The aim of this Section is to prove the following:
Theorem 4.3.
Let $U\subset L_{k}^{2n+1}$ and $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$. If $\Pi(\widetilde{\psi})$ $\alpha_{k}$-displaces $U$ then $\lceil C_{k}(U)\rceil_{\frac{2\pi}{k}}\leq c_{k}(\widetilde{\psi}^{-1})+\lceil c_{k}(\widetilde{\psi})\rceil_{\frac{2\pi}{k}}$. If moreover $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{S}^{1}}(L_{k}^{2n+1},\xi_{k})$ then $C_{k}(U)\leq c_{k}(\widetilde{\psi}^{-1})+c_{k}(\widetilde{\psi}).$
Remark 4.4.
  1. (1)
    The terminoly "energy-capacity inequality" is borrowed from symplectic geometry where a similar inequality was first described by Hofer and Zehnder while working with the Hofer-Zehnder capacity and the Hofer metric in the standard symplectic Euclidean space [25, Section 5] (see also [36, Prop. 4.12]). The proof of Theorem 4.3 is based on an adaptation of their proof.
  2. (2)
    One can use [4], (see for example the proof of [4, Prop 5.3] combined with the Poincaré duality of [4, Thm 1.1]), to show moreover that for every $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
    $0\leq\lceil c_{k}(\widetilde{\psi})\rceil_{\frac{2\pi}{k}}+\lfloor c_{k}(\widetilde{\psi}^{-1})\rfloor_{\frac{2\pi}{k}}\leq 2\pi+\frac{2\pi}{k}.$
    where $\lfloor x\rfloor_{\frac{2\pi}{k}}:=\frac{2\pi}{k}\left\lfloor x\cdot\frac{k}{2\pi}\right\rfloor$ is the biggest multiple of $\frac{2\pi}{k}$ smaller or equal to $x$ for any $x\in\mathbb{R}$. In particular if $U\subset L_{k}^{2n+1}$ is $\alpha_{k}$-displaceable then $\lceil C_{k}(U)\rceil_{\frac{2\pi}{k}}\leq 2\pi+\frac{4\pi}{k}$.
Theorem 4.3 will be a consequence of the next Proposition that can be stated for a general closed contact manifold $(M,\ker\alpha)$ admitting a $\alpha$-spectral selector. For any non-empty open set $U\subset M$, as before, $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi)\subset\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ denotes the subgroup of elements that can be represented by contact isotopies supported in $U$.
Proposition 4.5.
Let $U\subset M$ be a non-empty open set of a closed contact manifold $(M,\ker\alpha)$ and $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}^{c}(U,\ker\alpha)$. If $c:\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)\to\mathbb{R}$ is an $\alpha$-spectral selector and if $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)$ is such that $\Pi(\widetilde{\psi})$ $\alpha$-displaces $U\subset M$ then
$c(\widetilde{\psi}\cdot\widetilde{\varphi})=c(\widetilde{\psi}).$
Proposition 4.5 is a consequence of the following:
Lemma 4.6.
Let $\widetilde{\psi},\widetilde{\varphi}\in\mathrm{Cont}_{0}(M,\xi)$. If $\Pi(\widetilde{\psi})$ $\alpha$-displaces the support of $\Pi(\widetilde{\varphi})$ then
$\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})\subset\mathrm{Spec}^{\alpha}(\widetilde{\psi}).$
Proof of Lemma 4.6.
Let $t\in\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})$ and $x\in M$ so that $\phi_{\alpha}^{-t}\circ(\psi\circ\phi)$ where $\psi:=\Pi(\widetilde{\psi})$ and $\phi:=\Pi(\widetilde{\varphi})$. In particular $\psi(\phi(x))=\phi_{\alpha}^{t}(x)$. Suppose by contradiction that $x$ lies in the support of $\phi$. Then $\phi(x)\in\mathrm{Supp}(\phi)\subset\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))$ and $\phi_{\alpha}^{t}(x)\in\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))$. But this contradicts the fact that $\psi(\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))\cap\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))=\emptyset$. Therefore $x$ is not in the support of $\phi$. Thus $x$ is a discriminant point of $\phi_{\alpha}^{-t}\circ\psi$ and $t\in\mathrm{Spec}^{\alpha}(\widetilde{\psi})$. In particular $\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})\subset\mathrm{Spec}^{\alpha}(\widetilde{\psi})$. ∎
Proof of Proposition 4.5.
Let $(\varphi_{t})_{t\in[0,1]}$ be a contact isotopy supported in $U$ representing $\widetilde{\varphi}$ and $\widetilde{\varphi}_{s}:=[(\varphi_{ts})]$ for all $s\in[0,1]$. The map $\gamma:[0,1]\to\mathbb{R}$, $s\mapsto c(\widetilde{\psi}\cdot\widetilde{\varphi}_{s})$ is continuous (by Definition 2.1) and takes value in $\mathrm{Spec}^{\alpha}(\widetilde{\psi})$ thanks to Lemma 4.6. Since $\mathrm{Spec}^{\alpha}(\widetilde{\psi})$ is nowhere dense (see [1, Lemma 2.11]) we conclude that $\gamma$ is constant and in particular $\gamma(0)=c(\widetilde{\psi})=\gamma(1)=c(\widetilde{\psi}\cdot\widetilde{\varphi})$. ∎
We are now ready to prove Theorem 4.3:
Proof of Theorem 4.3.
Consider $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})$ then
$\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}=\lceil c_{k}(\widetilde{\psi}^{-1}\cdot\widetilde{\psi}\cdot\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}\leq c_{k}(\widetilde{\psi}^{-1})+\lceil c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}=c_{k}(\widetilde{\psi}^{-1})+\lceil c_{k}(\widetilde{\psi})\rceil_{\frac{2\pi}{k}},$ (2)
where the last equality follows from Proposition 4.5. This proves the first statement. Moreover under the assumption that $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{S}^{1}}(L_{k}^{2n+1},\xi_{k})$ one can use Proposition 2.5 to show that the relations in (2) remain true without taking the $\frac{2\pi}{k}$-ceiling function which proves the second statement. ∎
5. Estimation of the contact capacity
Let us recall our conventions to describe Hamiltonian dynamics. $(W,\Omega)$ being a symplectic manifold, to any compactly supported function $H:[0,1]\times W\to\mathbb{R}$ we associate the path of symplectic vector fields $(X_{H}^{t})$ on $W$, i.e. the flow $(\psi_{H}^{t})$ of which is an isotopy of symplectomorphisms, defined by
$\Omega(X_{t},\cdot)=-\mathrm{d}H(t,\cdot)\quad\text{ for all }t\in[0,1].$ (3)
We say that $H$ is the Hamiltonian function that generates the isotopy $(\psi_{H}^{t})$ and denote by $\mathrm{Ham}_{c}(W,\omega)$ the set of diffeomorphisms arising as time $1$-map of such isotopies.
Similarly if $(M,\xi=\ker\alpha)$ is a closed contact manifold, with a fixed contact form $\alpha$, to any function $h:[0,1]\times M\to\mathbb{R}$ one associates a path of contact vector fields $(Y_{h}^{t})$, i.e. the flow $(\phi_{h}^{t})$ of which is a contact isotopy, defined by the two equations
$\left\{\begin{array}[]{ll}\alpha(Y_{h}^{t})=h(t,\cdot)\\ \mathrm{d}\alpha(Y_{h}^{t},\cdot)=\mathrm{d}h^{t}(R_{\alpha})\alpha-\mathrm{d}h^{t}\end{array}\right.$ (4)
where $R_{\alpha}$ denotes the Reeb vector field associated to $\alpha$ (the vector field defined by the above equations (4) for the function constantly equal to $1$). We say that $h$ is the $\alpha$-Hamiltonian function that generates the contact isotopy $(\phi_{h}^{t})$.
We leave the proof of the following Lemma to the reader.
Lemma 5.1.
Suppose that $M$ is a $\mathbb{S}^{1}$-principle bundle over a closed symplectic manifold $(W,\Omega)$, that $M$ carries a contact form $\alpha$ satisfying $\pi^{*}\Omega=\mathrm{d}\alpha$, where $\pi:M\to W=M/\mathbb{S}^{1}$ denotes the projection, and that moreover the $\alpha$-Reeb flow induces the $\mathbb{S}^{1}$-action. Then for any function $H:[0,1]\times W\to\mathbb{R}$ the $\alpha$-Hamiltonian function
$h:[0,1]\times M\to\mathbb{R}\quad\quad\quad(t,x)\mapsto H(t,\pi(x))$
generates a contact isotopy $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(M,\ker\alpha)$ which is $\mathbb{S}^{1}$-equivariant, i.e. $\phi_{h}^{t}\circ\phi_{\alpha}^{s}=\phi_{\alpha}^{s}\circ\phi_{h}^{t}$ for all $s\in\mathbb{R}$, $t\in[0,1]$. Moreover $\pi\circ\phi_{h}^{t}=\psi_{H}^{t}\circ\pi$ for all $t\in[0,1]$, where $(\psi_{H}^{t})\subset\mathrm{Ham}(W,\Omega)$ is the isotopy generated by $H$.
5.1. Estimation of the lower bound
The aim of this subsection is to prove the lower bound in Theorem 3.1 satisfied by $C_{k}$, for any integer $k\geq 2$ and $n\in\mathbb{N}_{>0}$, which we recall in the next Proposition.
Proposition 5.2.
Let $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ be a symplectic embedding where $R>0$. Then $C_{k}(B_{k}(r))\geq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
The proof of Proposition 5.2 will be an adaptation of the proof of [25, Lemma 3, Chapter 3] stating that the Hofer-Zehnder capacity of $(\mathbb{B}^{2n}(r),\omega_{0})$ is bounded from below by $\pi r^{2}$.
Proof.
Step 1 : Construction in $(\mathbb{B}^{2n}(r),\omega_{0})$
Consider $0<\varepsilon<\pi r^{2}$ and a smooth function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ supported in $[0,r^{2})$ that satisfies moreover :
  1. (1)
    $f(t)=\pi r^{2}-\varepsilon$ for $t$ in a non-empty neighborhood of $\{0\}$
  2. (2)
    $-\pi<f^{\prime}(t)\leq 0$ for all $t\in\mathbb{R}_{\geq 0}$
  3. (3)
    $f^{\prime}(t)=0$ if and only if $f(t)\in\{\pi r^{2}-\varepsilon,0\}$
Figure 1: Graphs of different functions $f$ for different values of $\varepsilon$ with $0<\varepsilon_{3}<\varepsilon_{2}<\varepsilon_{1}$
[…]
$X_{H}(x)=J\nabla H(x)=2f^{\prime}(\|x\|^{2})Jx$
where $J$ is the standard complex structure of $\mathbb{R}^{2n}$555after identifying $\mathbb{R}^{2n}$ with $\mathbb{C}^{n}$ via $\Phi$ defined in the introduction $J:=\Phi^{-1}\circ i\circ\Phi$ corresponds to the multiplication by $i$ and $\nabla H$ denotes the gradient of $H$ with respect to the standard Euclidean scalar product. A direct computation shows that for all $t\in[0,1]$ and $x\in\mathbb{B}^{2n}(r)$
$\psi_{H}^{t}(x)=e^{2f^{\prime}(\|x\|^{2})Jt}x=\cos(2f^{\prime}(\|x\|^{2})t)x+\sin(2f^{\prime}(\|x\|^{2})t)Jx,$ (5)
where $(\psi_{H}^{t})\subset\mathrm{Ham}_{c}(\mathbb{B}^{2n}(r))$ denotes the flow generated by $H$.
Note that since $2f^{\prime}([0,+\infty[)\subset(-2\pi,0]$ and $f^{\prime}(0)=0$, the orbit $t\mapsto\psi_{H}^{t}(x)$ of a point $x\in\mathbb{B}^{2n}(r)$ is $T$-periodic for some $T\in(0,1]$, i.e. $\psi_{H}^{T}(x)=x$, if and only if $f^{\prime}(\|x\|^{2})=0$. In other words the orbit of $x$ is $T$-periodic, with $T\in(0,1]$, if and only if $f(\|x\|^{2})=H(x)\in\{\pi r^{2}-\varepsilon,0\}$ and $\mathrm{d}_{x}H=0$.
Step 2 : Import the construction to $(\mathbb{C}\mathrm{P}^{n},\omega)$ Consider now on $(\mathbb{C}\mathrm{P}^{n},\omega)$ the Hamiltonian function
$G:\mathbb{C}\mathrm{P}^{n}\to\mathbb{R}\quad\quad x\mapsto\left\{\begin{array}[]{ll}H(\Psi^{-1}(x))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(r))\\ 0&\text{ otherwise}\end{array}.\right.$
A direct computation shows that the isotopy it generates $(\psi_{G}^{t})\subset\mathrm{Ham}(\mathbb{C}\mathrm{P}^{n},\omega)$ satisfies :
$\psi_{G}^{t}:x\mapsto\left\{\begin{array}[]{ll}\Psi(\psi_{H}^{t}(\Psi^{-1}(x)))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(r))\\ x&\text{ otherwise}\end{array}.\right.$
In particular, thanks to the first step we get that the orbit $t\mapsto\psi_{G}^{t}(x)$ of a point $x\in\mathbb{C}\mathrm{P}^{n}$ is $T$-periodic, with $T\in(0,1]$, if and only if $G(x)\in\{\pi r^{2}-\varepsilon,0\}$ and $\mathrm{d}_{x}G=0$.
Step 3 : Import the construction to $(L_{k}^{2n+1},\xi_{k})$ Consider on $(L_{k}^{2n+1},\ker\alpha_{k})$ the $\alpha_{k}$-Hamiltonian function
$h:L_{k}^{2n+1}\to\mathbb{R},\quad\quad x\mapsto G(\pi_{k}(x)),$
and denote by $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ the flow of contactomorphisms it generates. It is clear that its lift $(\overline{\phi_{h}^{t}})\subset\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\ker\alpha)$, starting at the identity, is generated by the $\alpha$-contact Hamiltonian $\overline{h}:\mathbb{S}^{2n+1}\to\mathbb{R}$, $x\mapsto G(\pi_{1}(x))$, where for simplicity $\alpha:=\alpha_{1}$. We claim that: if $x\in\mathbb{S}^{2n+1}$ is a $\alpha$-translated point of $\overline{\phi_{h}^{T}}$ for some $T\in(0,1]$ then
there is $\text{there is }a\in\{\pi r^{2}-\varepsilon,0\}\text{ so that }\overline{\phi_{h}^{t}}(x)=\phi_{\alpha}^{at}(x)\text{ for all }t\in[0,1].$ (6)
To prove this claim, fix $x\in\mathbb{S}^{2n+1}$ and $T\in(0,1]$ so that $x$ is a $\alpha$-translated of $\overline{\phi_{h}^{T}}$. By Lemma 5.1 $x$ is a $\alpha$-translated point of $\overline{\phi_{h}^{T}}$ implies that $\pi_{1}(x)$ is a fixed point of $\psi_{G}^{T}$. This implies, by the previous step, that $G(\pi_{1}(x))\in\{\pi r^{2}-\varepsilon,0\}$ and $\mathrm{d}_{\pi_{1}(x)}G=0$. Therefore $\mathrm{d}_{x}\overline{h}=0$ and $\overline{h}(x)\in\{\pi r^{2}-\varepsilon,0\}$. Since $\overline{h}\circ\phi_{\alpha}^{t}=\overline{h}$ for all $t\in\mathbb{R}$ we also get
$\mathrm{d}\overline{h}(R_{\alpha})=0,\quad\mathrm{d}_{\phi_{\alpha}^{t}(x)}\overline{h}=0,\quad\text{ and }\quad\overline{h}(\phi_{\alpha}^{t}(x))=\overline{h}(x).$ (7)
The last three equations (7) together with equation (4) imply that the contact vector field $Y$ associated to the $\alpha$-Hamiltonian function $\overline{h}$ satisfies for all $t\in\mathbb{R}$
$Y(\phi_{\alpha}^{t}(x))=\overline{h}(x)R_{\alpha}(\phi_{\alpha}^{t}(x)).$
Therefore the map $[0,1]\to\mathbb{S}^{2n+1}$, $t\mapsto\phi_{\alpha}^{\overline{h}(x)t}(x)$ solves the differential equation $\gamma^{\prime}(t)=Y(\gamma(t))$ with initial condition $\gamma(0)=x$. By unicity of solutions of such equations $\overline{\phi_{h}^{t}}(x)=\phi_{\alpha}^{\overline{h}(x)t}(x)$ for all $t\in[0,1]$ and the claim (6) follows. In particular we have proved that for all $T\in(0,1]$
$\mathrm{Spec}^{\alpha}(\overline{\phi_{h}^{T}})=\left\{T(\pi r^{2}-\varepsilon)+2\pi m\ |\ m\in\mathbb{Z}\right\}\bigcup\left\{2\pi m\ |\ m\in\mathbb{Z}\right\}.$ (8)
Step 4 : Computation of the selector and conclusion Let us denote by $\widetilde{\phi_{h}^{T}}:=[(\phi_{h}^{tT})]\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ for all $T\in[0,1]$. Proposition 2.3 ensures that $c_{k}(\widetilde{\phi_{h}^{0}})=0$ and that the map $[0,1]\to\mathbb{R},\ T\mapsto c_{k}(\widetilde{\phi_{h}^{T}})\in\overline{\mathrm{Spec}^{\alpha_{k}}}(\widetilde{\phi_{h}^{T}})=\mathrm{Spec}^{\alpha}(\overline{\phi_{h}^{T}})$ is continuous. Therefore there is $a\in\{T(\pi r^{2}-\varepsilon),0\}$ so that $c_{k}(\widetilde{\phi_{h}^{T}})=Ta$ for all $T\in[0,\delta]$ where $\delta:=\min\{1,\frac{2\pi}{\pi r^{2}-\varepsilon}\}$.
Figure 2: The graphs of $f_{j}:T\mapsto T(\pi r^{2}-\varepsilon)+2\pi j$ and $g_{j}:T\mapsto 2\pi j$ for $j\in\{-1,0,1\}$
Since $\mathrm{id}\preceq\widetilde{\phi_{h}^{T}}$ and $\widetilde{\phi_{h}^{T}}\neq\mathrm{id}$, for $T>0$, we use Proposition 2.3 again to get that $c_{k}(\widetilde{\phi_{H}^{T}})>0$ and thus
$c_{k}(\widetilde{\phi_{h}^{T}})=T(\pi r^{2}-\varepsilon)\text{ for all }T\in[0,\delta].$ (9)
The relation (9) actually holds for all $T\in[0,1]$ since by [26, Thm 1] the radius $r\leq\sqrt{2}$ and so $\delta=1$ (see also Remark 1.2). The $\alpha_{k}$-Hamiltonian function $h$ being supported in $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n},(r)))$ we get by definition that
$C_{k}(B_{k}(r))\geq c_{k}(\widetilde{\phi_{h}^{1}})=\pi r^{2}-\varepsilon.$
Since $\varepsilon>0$ can be choosen as small as we want this finishes the proof. ∎
5.2. Estimation of the upper bound
The aim of this subsection is to prove the upper bound in Theorem 3.1, for all $n\in\mathbb{N}_{>0}$ and integer $k\geq 2$, which we recall in the following Proposition.
Proposition 5.3.
Consider $\Psi:(\mathbb{B}^{2n}(\lambda R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ a symplectic embedding where $R>0$ and $\lambda>\sqrt{3}$. Then for all $r\in(0,R]$
$C_{k}(B_{k}(r))\leq\pi r^{2}$
where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
Let us first consider $(\mathbb{R}^{2}=\mathbb{R}^{2}\times\{0\},dx_{1}\wedge dy_{1})$ the symplectic submanifold of $(\mathbb{R}^{2n},\sum\limits_{i=1}^{n}dx_{i}\wedge dy_{i})$. We recall a Lemma from [24, Lemma 4.2].
Lemma 5.4 ([24]).
For every $\delta>0$ there exists an Hamiltonian function $H:\mathbb{R}^{2}\to\mathbb{R}_{\geq 0}$ supported in $\mathbb{B}^{2}(\sqrt{2+2\delta})$ so that $\max H\leq(\pi+\delta)$ and $\psi_{H}^{1}(\mathbb{B}^{2}(1))\bigcap\mathbb{B}^{2}(1)=\emptyset$.
As a direct corollary we have.
Corollary 5.5.
For every $r,\delta>0$ there exists an Hamitlonian function $H:\mathbb{R}^{2}\to\mathbb{R}_{\geq 0}$ supported in $\mathbb{B}^{2}(r\sqrt{2+2\delta})$ so that $\max H\leq r^{2}(\pi+\delta)$ and $\psi_{H}^{1}(\mathbb{B}^{2}(r))\bigcap\mathbb{B}^{2}(r)=\emptyset$.
Proof.
Let $H$ be a Hamiltonian function that satisfies the requirements of Lemma 5.4 with respect to $\delta>0$ and consider the diffeomorphism $\psi_{r}:\mathbb{R}^{2}\to\mathbb{R}^{2},\ x\mapsto rx$. A direct computation shows that the isotopy $(\psi_{r}\circ\psi_{t}\circ\psi_{r}^{-1})$ is contained in $\mathrm{Ham}_{c}(\mathbb{R}^{2},\omega_{0})$ and is generated by the Hamiltonian function $H_{r}:=r^{2}H\circ(\psi_{r})^{-1}$ whose support is equal to $(\psi_{r})(\operatorname{supp}(H))$. Therefore $H_{r}$ satisfies the requirements of Corollary 5.5 with respect to $\delta>0$ and $r>0$. ∎
We are now ready to prove Proposition 5.3.
Proof of Proposition 5.3.
Step 1 : Construction in $(\mathbb{B}^{2n}(\lambda R),\omega_{0})$
Let $r\in(0,R]$, $\delta>0$ be small enough so that $\sqrt{3+2\delta}<\lambda$ and $H:\mathbb{R}^{2}\to\mathbb{R}_{\geq 0}$ be an Hamiltonian function supported in $\mathbb{B}^{2}(r\sqrt{2+2\delta})$ so that $\max H\leq r^{2}(\pi+\delta)$ and $\psi_{H}^{1}(\mathbb{B}^{2}(r))\bigcap\mathbb{B}^{2}(r)=\emptyset$ whose existence is guaranteed by Corollary 5.5. The function $K:\mathbb{R}^{2n}\to\mathbb{R}_{\geq 0},((x_{1},y_{1}),\cdots,(x_{n},y_{n}))\mapsto H(x_{1},y_{1})$, even if non compactly supported, generates a vector field $X_{k}$, i.e. $\omega_{0}(X_{k},\cdot)=-dK$, which is complete, and whose flow $(\psi_{K}^{t})\subset\mathrm{Ham}(\mathbb{R}^{2n},\omega_{0})$ satisfies $\psi_{K}^{t}=\psi_{H}^{t}\times\mathrm{id}_{\mathbb{R}^{2(n-1)}}$ for all $t\in[0,1]$. Therefore
$\psi_{K}^{1}\left(\mathbb{B}^{2n}(r)\right)\bigcap\mathbb{B}^{2n}(r)\subset\left(\psi_{H}^{1}\left(\mathbb{B}^{2}(r)\right)\times\mathbb{B}^{2(n-1)}(r)\right)\bigcap\mathbb{B}^{2}(r)\times\mathbb{B}^{2(n-1)}(r)=\emptyset.$
Moreover $\underset{t\in[0,1]}{\bigcup}\psi_{K}^{t}(\mathbb{B}^{2n}(r))$ is contained in $\mathbb{B}^{2n}(r\sqrt{3+2\delta})$ since
$\psi_{K}^{t}(\mathbb{B}^{2n}(r))\subset\psi_{H}^{t}(\mathbb{B}^{2}(r))\times\mathbb{B}^{2(n-1)}(r)\subset\mathbb{B}^{2}(r\sqrt{2+2\delta})\times\mathbb{B}^{2(n-1)}(r)$
and the latter set is clearly contained inside $\mathbb{B}^{2n}(r\sqrt{3+2\delta})$. Given that $\sqrt{3+2\delta}<\lambda$ one can construct $\rho:\mathbb{R}^{2n}\to[0,1]$ a smooth cutoff function supported in $\mathbb{B}^{2n}(\lambda R)$ which is constantly equal to $1$ on $\mathbb{B}^{2n}(r\sqrt{3+2\delta})$ so that the Hamiltonian function $\widetilde{K}:=\rho K$ generates a path of Hamiltonian symplectomorphisms $(\psi_{t})$ satisfying
  1. (1)
    $\widetilde{K}$ and $(\psi_{t})$ are both supported in $\mathbb{B}^{2n}(\lambda R)$
  2. (2)
    $\widetilde{K}\geq 0$ and $\max\widetilde{K}\leq r^{2}(\pi+\delta)$
  3. (3)
    $\psi_{1}\left(\mathbb{B}^{2n}(r)\right)\cap\mathbb{B}^{2n}(r)=\emptyset$.
Step 2 : Import the construction to $(\mathbb{C}\mathrm{P}^{n},\omega)$ A direct computation shows that the Hamiltonian function
$G:\mathbb{C}\mathrm{P}^{n}\to\mathbb{R}_{\geq 0},\quad\quad x\mapsto\left\{\begin{array}[]{ll}\widetilde{K}(\Psi^{-1}(x))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(\lambda R))\\ 0&\text{ otherwise}\end{array}\right.$
on $(\mathbb{C}\mathrm{P}^{n},\omega)$, generates an isotopy $(\psi_{G}^{t})\subset\mathrm{Ham}(\mathbb{C}\mathrm{P}^{n},\omega)$ that satisfies
$[0,1]\times\mathbb{C}\mathrm{P}^{n}\to\mathbb{C}\mathrm{P}^{n},\quad\quad(t,x)\mapsto\left\{\begin{array}[]{ll}\Psi(\psi_{t}(\Psi^{-1}(x)))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(\lambda R))\\ x&\text{ otherwise.}\end{array}\right.$
We therefore clearly have:
  1. (1)
    $G\geq 0$ and $\max G\leq r^{2}(\pi+\delta)$
  2. (2)
    $\psi_{G}^{1}(B(r))\cap B(r)=\emptyset$, where $B(r):=\Psi(\mathbb{B}^{2n}(r))$.
Step 3 : Lift the construction to $(L_{k}^{2n+1},\xi_{k})$ We lift this construction to $(L_{k}^{2n+1},\ker\alpha_{k})$, thanks to Lemma 5.1. More precisely the $\alpha_{k}$-Hamiltonian function $h:L_{k}^{2n+1}\to\mathbb{R}_{\geq 0}$, $x\mapsto G(\pi_{k}(x))$ generates a flow of $\mathbb{S}^{1}$-equivariant contactomorphisms $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ satisfying :
  1. (1)
    $h\geq 0$ and $\max h\leq r^{2}(\pi+\delta)$
  2. (2)
    $\phi_{h}^{1}(B_{k}(r))\cap B_{k}(r)=\emptyset$, where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
Indeed, if we suppose by contradiction that $\phi_{h}^{1}(B_{k}(r))\cap B_{k}(r)\neq\emptyset$ then we get the following absurdity
$\emptyset\neq\pi_{k}\left(\phi_{h}^{1}(B_{k}(r))\cap B_{k}(r)\right)\subset\pi_{k}\left(\phi_{h}^{1}(B_{k}(r))\right)\cap\pi_{k}\left(B_{k}(r)\right)=\psi_{G}^{1}(B(r))\cap B(r)=\emptyset.$
Step 4 : Conclusion Therefore thanks to Theorem 4.3
$c_{k}(B_{k}(r))\leq c_{k}(\widetilde{\phi}^{-1})+c_{k}(\widetilde{\phi})$
where $\widetilde{\phi}:=[(\phi_{t})]\in\widetilde{\mathrm{Cont}_{0}}(L_{K}^{2n+1},\xi_{k})$. Finally, since the $\alpha_{k}$-Hamiltonian function $L_{k}^{2n+1}\to\mathbb{R}_{\leq 0},\ x\mapsto-h(x)$, generates the contact isotopy $(\phi_{t})^{-1}$ which represents $\widetilde{\phi}^{-1}$, the last point of Proposition 2.3 implies that
$c_{k}(\widetilde{\phi})+c_{k}(\widetilde{\phi}^{-1})\leq\int_{0}^{1}\max h\mathrm{d}t-\int_{0}^{1}\min h\mathrm{d}t\leq r^{2}(\pi+\delta).$
Letting $\delta$ goes to 00 finishes the proof. ∎
6. Possible generalization to orderable prequantizations
6.1. Orderability and pseudo-spectral selectors
Let $(M,\xi)$ be a closed coorientable contact manifold and from now on we fix a coorientation of $\xi$, i.e. a vector field $X\in\chi(M)$ so that $X(M)\subset TM\setminus\xi$. We say that a contact form $\alpha$ supports $\xi$ if $\ker\alpha=\xi$ and $\alpha(X)>0$. A contact isotopy $(\phi_{t})$ is said to be non-negative if for some, and thus all, contact form $\alpha$ supporting $\xi$ we have $\alpha(\frac{d}{dt}\phi_{t}(x))\geq 0$ for all $(t,x)\in[0,1]\times M$. Recall from [15] that $(M,\xi)$ is said to be orderable if the only non-negative contact isotopy $(\phi_{t})$ representing $\mathrm{id}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ is the constant isotopy. Under the assumption that $(M,\xi)$ is orderable, in [3, 5, 30] the authors show that for any contact form $\alpha$ supporting $\xi$ the map $c_{+}^{\alpha}:\widetilde{\mathrm{Cont}_{0}}(M,\xi)\to\mathbb{R}$ defined by
[…]
is a well defined $C^{1}$-continuous map, where $\widetilde{\varphi}\preceq\widetilde{\psi}$ if and only if $\widetilde{\psi}\cdot\widetilde{\varphi}^{-1}$ can be represented by a non-negative contact isotopy $(\phi_{t})$. Moreover it is shown in [3, 30] that this map shares a lot of properties with the $\alpha$-spectral selector of Proposition 2.3 constructed in [4]. Namely for all $\widetilde{\varphi},\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)$
  1. (1)
    $c_{+}^{\alpha}(\mathrm{id})=0$
  2. (2)
    $\widetilde{\varphi}\preceq\widetilde{\psi}$ implies that $c^{\alpha}_{+}(\widetilde{\varphi})\leq c^{\alpha}_{+}(\widetilde{\psi})$
  3. (3)
    $\mathrm{id}\preceq\widetilde{\varphi}$ and $\widetilde{\varphi}\neq\mathrm{id}$ implies that $c^{\alpha}_{+}(\widetilde{\varphi})>0$
  4. (4)
    $c_{+}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})\leq c_{+}^{\alpha}(\widetilde{\psi})+c_{+}^{\alpha}(\widetilde{\varphi})$
  5. (5)
    for any contact isotopy $(\phi_{t})$ representing $\widetilde{\varphi}$
    $\int_{0}^{1}\min\alpha\left(\frac{d}{dt}\phi_{t}\right)\ \mathrm{d}t\leq c_{+}^{\alpha}(\widetilde{\varphi})\leq\int_{0}^{1}\max\alpha\left(\frac{d}{dt}\phi_{t}\right)\ \mathrm{d}t.$
As in Section 3 defining a map $C_{\alpha}:\{\text{non-empty open set of }M\}\to\mathbb{R}_{>0}\cup\{+\infty\}$
$U\mapsto C_{\alpha}(U):=\sup\{c_{+}^{\alpha}(\widetilde{\varphi})\ |\ \widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi)\},$
we get that
if $\text{if }U\subset V\text{ then }C_{\alpha}(U)\leq C_{\alpha}(V).$ (10)
Without making more assumption on the contact manifold $(M,\xi)$ and the map $c_{+}^{\alpha}$ one cannot say much more about $C_{\alpha}$. However in [3] the authors made the following conjecture
Conjecture 6.1.
If $(M,\xi)$ is a closed orderable contact manifold then $c_{+}^{\alpha}$ is an $\alpha$-spectral selector for any contact form $\alpha$ supporting $\xi$.
The next Lemma gives an important feature of $C_{\alpha}$ under the assumptions that the conjecture 6.1 holds and that the $\alpha$-Reeb flow is periodic.
Lemma 6.2.
Let $(M,\xi)$ be a orderable closed contact manifold and $\alpha$ be a contact form supporting $\xi$ whose Reeb flow is periodic of minimal period $T>0$. If $c:\widetilde{\mathrm{Cont}_{0}}(M,\xi)\to\mathbb{R}$ is a $\alpha$-spectral selector which is monotone, i.e. $\widetilde{\varphi}_{1}\preceq\widetilde{\varphi}_{2}$ implies that $c(\widetilde{\varphi}_{1})\leq c(\widetilde{\varphi}_{2})$, then
$\lceil c(\widetilde{\psi}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}^{-1})\rceil_{T}=\lceil c(\widetilde{\varphi})\rceil_{T}\quad\quad\text{ for all }\widetilde{\varphi},\ \widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi).$
In particular for all non-empty open set $U$ in $M$ and $\psi\in\mathrm{Cont}_{0}(M,\xi)$
$\lceil C(U)\rceil_{T}=\lceil C(\psi(U))\rceil_{T},$
where $C(V):=\sup\{c(\widetilde{\varphi})\ |\ \widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(V,\xi)\}$ for all non-empty open sets $V\subset M$.
Proof.
Consider $(\psi_{t})$ a path representing $\widetilde{\psi}$ and denote by $\widetilde{\psi}_{s}:=[(\psi_{st})]$ for all $s\in[0,1]$. Let us first suppose that $\Pi(\widetilde{\varphi})$ does not have discriminant point. Thus $\Pi(\widetilde{\psi}_{s}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{s}^{-1})$ does not have discriminant point and so $c(\widetilde{\psi}_{s}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{s}^{-1})\notin T\mathbb{Z}$ for all $s\in[0,1]$. We then conclude using the continuity of $s\mapsto c(\widetilde{\psi}_{s}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{s}^{-1})$. We treat the general case using the fact that $\mathrm{Spec}^{\alpha}(\widetilde{\varphi})$ is nowhere dense (see [1, Lemma 2.11]) and therefore there exists a sequence $(\varepsilon_{n})_{n\in\mathbb{N}}$ of positive real numbers going to 00 so that the sequence $\Pi(\widetilde{\phi_{\alpha}^{-\varepsilon_{n}}}\cdot\widetilde{\varphi})$ does not have any discriminant for all $n\in\mathbb{N}$. Since $\widetilde{\phi_{\alpha}^{-\varepsilon_{n}}}\cdot\widetilde{\varphi}\preceq\widetilde{\varphi}$ and the sequence $(\widetilde{\phi_{\alpha}^{-\varepsilon_{n}}}\cdot\widetilde{\varphi})$ converges to $\widetilde{\varphi}$ in the $C^{1}$-topology we conclude using the previous case and the monotonicity of the $\alpha$-selector. ∎
6.2. Prequantization
Let $(W,\omega)$ be a closed symplectic manifold so that the natural morphism $H^{2}(W,\mathbb{Z})\to H^{2}(W,\mathbb{R})$ is without torsion, i.e. it is injective, $[\frac{\omega}{2\pi}]$ lies in its image and $[\omega],H_{2}(W,\mathbb{Z})\right>=2\pi\mathbb{Z}$. Then there exists a unique, up to $\mathbb{S}^{1}$-principal bundle isomorphism, $\mathbb{S}^{1}$-principal bundle $\pi_{1}:M_{1}\to W$ whose Euler class $eu(\pi_{1}:M_{1}\to W)=-[\frac{\omega}{2\pi}]$, where $\mathbb{S}^{1}=\{e^{i\theta}\ |\ \theta\in[0,2\pi]$}. Moreover $M_{1}$ carries a contact form $\alpha_{1}$ that satisfies $\pi_{1}^{*}\omega=\mathrm{d}\alpha_{1}$, the Reeb flow of which is Zoll of minimal period $2\pi$ and induces the $\mathbb{S}^{1}$ action on $M_{1}$. Finally, for any integer $k\geq 1$, seeing $\mathbb{Z}_{k}$ as the subgroup of $k$th roots of unity in $\mathbb{S}^{1}$ we have a bundle $\pi_{k}:M_{k}\to W$, where $M_{k}:=M_{1}/\mathbb{Z}_{k}$, with Euler class $-k[\frac{\omega}{2\pi}]$. This manifold $M_{k}$ carries a contact form $\alpha_{k}$ defined by $(\pi^{k}_{1})^{*}\alpha_{k}=\alpha_{1}$ where $\pi^{k}_{1}:M_{1}\to M_{k}$ is the covering map and it satisfies $\mathrm{d}\alpha_{k}=\pi_{k}^{*}\omega$. Moreover the Reeb flow of $\alpha_{k}$ is Zoll of minimal period $\frac{2\pi}{k}$ (see for instance the introduction of [34] and [19, Thm 7.2.4]). For all $k\in\mathbb{N}_{>0}$ we denote by $\xi_{k}:=\ker\alpha_{k}$ the contact distribution on $M_{k}$ and fix the coorientation given by the Reeb vector field.
Lemma 6.3.
Suppose there exists $k\in\mathbb{N}_{>0}$ such that $(M_{k},\xi_{k})$ is orderable. Then $(M_{Nk},\xi_{kN})$ is orderable for any $N\in\mathbb{N}_{>0}$. Moreover for any non-empty open set $U\subset W$
$C_{\alpha_{Nk}}(\pi_{Nk}^{-1}(U))\geq C_{\alpha_{k}}(\pi_{k}^{-1}(U)).$ (11)
Proof.
We leave the proof of the first statement to the reader. Let us thus prove the inequality (11). First note that $c_{+}^{\alpha_{Nk}}(\widetilde{\varphi})\geq c_{+}^{\alpha_{k}}(\mathcal{L}(\widetilde{\varphi}))$ for all $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M_{Nk},\xi_{Nk})$ where $\mathcal{L}:\widetilde{\mathrm{Cont}_{0}}(M_{Nk},\xi_{Nk})\to\widetilde{\mathrm{Cont}_{0}}(M_{k},\xi_{k})$ denotes the lifting map. Then remark that for all $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{k}^{-1}(U),\xi_{k})$ there exists $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{Nk}^{-1}(U),\xi_{Nk})$ whose lift $\mathcal{L}(\widetilde{\psi})$ lies in $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{k}^{-1}(U),\xi_{k})$ and is such that $\widetilde{\varphi}\preceq\mathcal{L}(\widetilde{\psi})$. ∎
We have therefore the next Proposition.
Proposition 6.4.
Let $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(W,\omega)$ be a symplectic embedding and suppose that $(M_{k},\ker\alpha_{k})$ is orderable for some integer $k$. If Conjecture 6.1 holds then there exists $N_{0}$ and $a_{0}\in(0,R]$ so that for all $N>N_{0}$ the open set $B_{Nk}(R)$ cannot be contactly squeezed inside $B_{Nk}(a)$ for all $a\in(0,a_{0}]$ where $B_{Nk}(r):=\pi_{Nk}^{-1}(\Psi(\mathbb{B}^{2n}(R))$ are seen as subset of the contact manifold $(M_{Nk},\ker\alpha_{Nk})$.
Proof.
Put $\delta:=C_{\alpha_{k}}(B_{k}(R))>0$. For $N_{0}\in\mathbb{N}_{>0}$ big enough $\delta>\frac{2\pi}{kN_{0}}$. Since, for any $N\in\mathbb{N}_{>0}$, $(M_{Nk},\xi_{Nk})$ is orderable and since we have assumed that Conjecture 6.1 holds it implies that one can apply the capacity-energy inequality of Section 4 and the computation in 5.2 to get that
$C_{\alpha_{Nk}}(B_{Nk}(a))\leq\pi a^{2}\leq\frac{\pi}{kN_{0}}$
for all $a\in(0,a_{0}]$, where $a_{0}:=\min\{\frac{R}{\sqrt{3}},(kN_{0})^{-\frac{1}{2}}\}$. In addition to that, thanks to Lemma 6.3
$\frac{2\pi}{kN_{0}}<C_{\alpha_{k}}(B_{k}(R))\leq C_{\alpha_{Nk}}(B_{Nk}(R)).$
Suppose by contradiction that there exists $\phi\in\mathrm{Cont}_{0}(M_{Nk},\xi_{k})$ so that $\phi(B_{Nk}(R))\subset B_{Nk}(a)$ for some integer $N>N_{0}$. Then by (10)
$C_{\alpha_{Nk}}(\phi(B_{Nk}(R)))\leq C_{\alpha_{Nk}}(B_{Nk}(a))\leq\frac{\pi}{kN_{0}}.$
Moreover, under the assumption that Conjecture 6.1 holds, one can use Lemma 6.2 to get that
$\lceil C_{\alpha_{Nk}}(\phi(B_{Nk}(R)))\rceil_{\frac{2\pi}{Nk}}=\lceil C_{\alpha_{Nk}}(B_{Nk}(R))\rceil_{\frac{2\pi}{Nk}}.$
We therefore get the following contradiction
$\frac{2\pi}{Nk}\left\lceil\frac{N}{N_{0}}\right\rceil=\left\lceil\frac{2\pi}{kN_{0}}\right\rceil_{\frac{2\pi}{Nk}}\leq\lceil C_{\alpha_{Nk}}(\phi(B_{Nk}(R)))\rceil_{\frac{2\pi}{Nk}}\leq\left\lceil\frac{\pi}{kN_{0}}\right\rceil_{\frac{2\pi}{Nk}}=\frac{2\pi}{Nk}\left\lceil\frac{1}{2}\cdot\frac{N}{N_{0}}\right\rceil.$