### v2

# Contact non-squeezing in various closed prequantizations

Abstract.
We describe some contact non-squeezing phenomena, first in lens spaces then in strongly orderable closed prequantizations in the sense of Liu [30]. To detect and quantify these non-squeezing phenomena we define and compute two different contact capacities. In the former case the contact capacity comes from the spectral selectors constructed by Allais, Sandon and the author by the means of generating functions and Givental’s non-linear Maslov index [4]. In the latter case the contact capacity comes from the order spectral selectors constructed by Allais and the author [3].
Key words and phrases:
Contact geometry, non-squeezing, lens spaces, generating functions, Givental’s non-linear Maslov index, spectral invariants, orderability, contactomorphisms group
2020 Mathematics Subject Classification:
53D10, 57R17, 58B20
1. Introduction
Gromov’s non-squeezing theorem [23] is one of the cornerstones, if not the cornestone, of symplectic geometry. Indeed, while it had been known for a long time that a symplectic manifold does not carry any local invariant, the existence of global invariants finer than the volume was not clear until Gromov’s work in [23] (see also [24, 13]). Originally stated for the standard symplectic Euclidean space, this theorem has then been generalized to all symplectic manifolds by Lalonde-McDuff [28].
For contact geometry, considered sometimes as the odd-dimensional analogue of symplectic geometry, the story is somewhat different. In this geometry, the volume is not an invariant anymore and, depending on the contact manifold considered, global invariants might not exist at all. For example, in the standard contact Euclidean space, which is the local model of contact geometry, any bounded domain can be contactly squeezed inside any arbitrary small neighborhood. Therefore, in a general contact manifold, non-squeezing phenomena cannot be expected at small scale. Nonetheless, 20 years after Gromov’s result Eliashberg, Kim and Polterovich [15] overcame this additional difficulty and proved a non-squeezing theorem at large scale in the prequantization of the standard symplectic Euclidean space. Unlike symplectic geometry, generalization of non-squeezing phenomena to other contact manifolds remain however rare. To the knowledge of the author, such phenomena have been detected only in prequantization of Liouville manifolds ([1, 5, 2, 36, 17, 9, 19]), some exotic contact spheres [38] and the unitary cotangent of the torus endowed with its canonical contact structure (see for example [14, Thm 2.4.1]).
In this paper we show that interesting non-squeezing phenomena happen also in lens spaces endowed with their standard contact structure and in strongly orderable closed prequantizations.
1.1. Non-squeezing in contact lens spaces
Let us fix $n\in\mathbb{N}_{>0}$ a positive integer. Endow the Euclidean space $\mathbb{R}^{2n+2}$ with the coordinate functions $((x_{1},y_{1}),\cdots,(x_{n+1},y_{n+1}))$ and let $\mathbb{S}^{2n+1}:=\left\{\sum\limits_{j=1}^{n+1}x_{j}^{2}+y_{j}^{2}=1\right\}$ be the standard Euclidean sphere of radius $1$ centered at 00. Identifying $\mathbb{R}^{2n+2}$ with $\mathbb{C}^{n+1}$ via $\Phi:((x_{1},y_{1}),\cdots,(x_{n+1},y_{n+1}))\mapsto(x_{1}+iy_{1},\cdots,x_{n+1}+iy_{n+1})$ we get an action of $\mathbb{S}^{1}:=\{e^{i\theta}\ |\ \theta\in\mathbb{R}\}\subset\mathbb{C}$ on $\mathbb{S}^{2n+1}\subset\mathbb{C}^{n+1}$ given by multiplication. For all integer $k\geq 1$, the lens space $L_{k}^{2n+1}$ is the quotient space $\mathbb{S}^{2n+1}/\mathbb{Z}_{k}$ where $\mathbb{Z}_{k}:=\mathbb{Z}/k\mathbb{Z}\simeq\{e^{\frac{2i\pi m}{k}}\ |\ m\in[0,k-1]\cap\mathbb{N}\}$ is identified with the $k$th roots of unity. It carries a contact distribution $\xi_{k}$ given by the kernel of the contact form $\alpha_{k}$ defined as the pullback, by the covering map $\mathbb{S}^{2n+1}\to L_{k}^{2n+1}$, of $\alpha_{1}:=i_{\mathbb{S}^{2n+1}}^{*}\left(\sum\limits_{j=1}^{n+1}(x_{j}dy_{j}-y_{j}dx_{j})\right)$ where $i_{\mathbb{S}^{2n+1}}:\mathbb{S}^{2n+1}\hookrightarrow\mathbb{R}^{2n+2}$ is the inclusion.
Finally we identify the complex projective space $\mathbb{C}\mathrm{P}^{n}$ with the quotient space $\mathbb{S}^{2n+1}/\mathbb{S}^{1}$ and denote by $\pi_{1}:\mathbb{S}^{2n+1}\to\mathbb{C}\mathrm{P}^{n}$ the projection. We consider on $\mathbb{C}\mathrm{P}^{n}$ the only symplectic form $\omega$ that satisfies $\pi_{1}^{*}\omega:=\mathrm{d}\alpha_{1}$. Note that $\omega=2\omega_{\mathrm{FS}}$, where $\omega_{\mathrm{FS}}$ denotes the standard Fubini-Study form, and that $\pi_{k}^{*}\omega=\mathrm{d}\alpha_{k}$, where $\pi_{k}:L_{k}^{2n+1}\to\mathbb{C}\mathrm{P}^{n}$ is the natural projection.
$M$ being a smooth manifold and $\xi$ a contact distribution on $M$, we say that a subset $U_{1}\subset M$ cannot be contactly squeezed inside another subset $U_{2}\subset M$ if for all contactomorphisms $\phi\in\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ isotopic to the identity through a compactly supported contact isotopy, $\phi(U_{1})$ does not lie inside $U_{2}$, i.e. $\phi(U_{1})\cap U_{2}\neq\phi(U_{1})$.
Denoting by $\mathbb{B}^{2n}(r):=\{((x_{1},y_{1}),\cdots,(x_{n},y_{n}))\in\mathbb{R}^{2n}\ |\ \sum\limits_{j=1}^{n}x_{j}^{2}+y_{j}^{2}<r^{2}\}$ the standard open Euclidean ball of radius $r>0$ in $\mathbb{R}^{2n}$ centered at 00 and by $\omega_{0}:=\sum\limits_{j=1}^{n}dx_{j}\wedge dy_{j}$ the standard symplectic form of $\mathbb{R}^{2n}$ we will prove the following:
Theorem 1.1.
Let $\Psi_{1}:(\mathbb{B}^{2n}(R_{1}),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ and $\Psi_{2}:(\mathbb{B}^{2n}(\lambda R_{2}),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ be two symplectic embeddings for some $R_{1},\ R_{2}>0$, $\lambda>\sqrt{3}$ and let $a_{1}\in(0,R_{1}],\ a_{2}\in(0,R_{2}]$ be two positive real numbers. For any integer $k\geq 2$, if there exists $j\in\mathbb{N}_{>0}$ such that
$\pi a_{1}^{2}>\frac{2\pi}{k}\cdot j\geq\pi a_{2}^{2},$
then $B^{1}_{k}(a_{1})$ cannot be contactly squeezed inside $B^{2}_{k}(a_{2})$, where by definition $B^{i}_{k}(a_{i}):=\pi_{k}^{-1}(\Psi_{i}(\mathbb{B}^{2n}(a_{i})))\subset(L_{k}^{2n+1},\xi_{k})$ for all $i\in\{1,2\}$.
Remark 1.2.
  1. (1)
    It might be that Theorem 1.1 remains true for some $\lambda\leq\sqrt{3}$. However, for the purpose of our proof, $\lambda$ has to be strictly greater than $1$ since, as we will see in Subsection 6.2, we need to displace $\mathbb{B}^{2n}(a_{2})$ from itself inside $\mathbb{B}^{2n}(\lambda R_{2})$.
  2. (2)
    These non-squeezing phenomena do not happen if we forget about the contact structure. Indeed it is easy to see that if $R_{1}=\lambda R_{2}=:R$ and $\Psi_{1}=\Psi_{2}=:\Psi$ then for all $r_{1},\ r_{2}\in(0,R]$, there exists a diffeomorphism of $L_{k}^{2n+1}$ which is isotopic to the the identity - through diffeomorphisms - and that sends $B_{k}(r_{1})$ inside $B_{k}(r_{2})$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$ for all $r\in(0,R]$.
  3. (3)
    In [27, Thm 1 and Prop 2.11] [^1] The conventions used in [27] are different: they consider the symplectic form $\widetilde{\omega}:=\frac{1}{2\pi}{\omega}$ and show that there exists an embedding of $(\widetilde{\mathbb{B}^{2n}}(1),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\widetilde{\omega})$, where $\widetilde{\mathbb{B}^{2n}}(a):=\mathbb{B}^{2n}(\sqrt{\frac{a}{\pi}})$ Karshon and Tolman construct a symplectic embedding $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$, for $R=\sqrt{2}$, and show that it is optimal in the sense that such symplectic embeddings do not exist for $R>\sqrt{2}$ (see also [31, Thm 1]). Considering this optimal embedding, one can see that non-squeezing phenomena appear for all lens spaces, even for the real projective space $(\mathbb{R}\mathrm{P}^{2n+1},\xi):=(L_{2}^{2n+1},\xi_{2})$, since $\pi\left(\sqrt{2}\right)^{2}>\frac{2\pi}{2}$.
We will prove Theorem 1.1 by constructing a contact capacity defined on open subsets of $(L_{k}^{2n+1},\xi_{k})$ and by estimating this contact capacity for the lift of Darboux balls. This contact capacity comes from the spectral selectors defined on $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ in [4] by the means of generating functions and Givental’s non-linear Maslov index (see Section 3).
1.2. Non-squeezing in strongly orderable prequantizations
Recall that if $(M,\xi=\ker\alpha)$ is a closed cooriented contact manifold then $(M\times M\times\mathbb{R},\Xi:=\ker(\beta:=\mathrm{pr}_{2}^{*}\alpha-e^{\theta}\mathrm{pr}_{1}^{*}\alpha))$, where $\mathrm{pr}_{i}:M\times M\times\mathbb{R}\to M,\ (x_{1},x_{2},\theta)\mapsto x_{i}$, is a cooriented contact manifold and the diagonal $\Delta:=\{(x,x,0)\ |\ x\in M\}$ is a Legendrian lying inside it. We will call the latter manifold the contact product of $(M,\xi)$222the contact structure on the contact product of $(M,\xi)$ does not depend on the choice of a contact form defining the contact distribution $\xi$ up to isomorphism with itself and will call the $1$-form $\beta$ the product contact form associated to $\alpha$. Using the terminology of [29, 7] we say that $(M,\ker\alpha)$ is strongly orderable if there does not exist any positive contractible loop of Legendrians based at $\Delta$. In particular, if $(M,\ker\alpha)$ is strongly orderable then it is orderable in the sense of [16], i.e. there does not exist any positive contractible loop of contactomorphisms in $\mathrm{Cont}_{0}(M,\ker\alpha)$. See Sections 2.2 and 7.1 for more details about these notions. Colin-Chantraine-Dimitroglou Rizell show in [7, Thm 1.19] that if $(M,\xi)$ is a closed hypertight contact manifold, i.e. there exists a contact form $\alpha$ supporting $\xi$ whose Reeb flow does not have any contractible Reeb orbit, then $(M,\xi)$ is strongly orderable. They also show that any closed contact manifold admitting a Liouville filling whose symplectic homology does not vanish is strongly orderable [7, Thm 1.18].
We say that a closed contact manifold $(M_{1},\xi_{1}^{\prime})$ is a prequantization over a closed symplectic manifold $(W,\omega^{\prime})$ if there exists a $\mathbb{T}^{1}$-principal bundle $\pi_{1}^{\prime}:M_{1}\to W$, where $\mathbb{T}^{1}=\mathbb{R}/2\pi\mathbb{Z}$, and a contact form $\alpha_{1}^{\prime}$ supporting $\xi_{1}^{\prime}$ satisfying ${\pi_{1}^{\prime}}^{*}\omega^{\prime}=\mathrm{d}\alpha^{\prime}_{1}$ and the Reeb flow of which, is Zoll333all its orbits are periodic with the same minimal period of minimal period $2\pi$, induces the $\mathbb{T}^{1}$-action.
Suppose that $(M_{1},\xi_{1}^{\prime})$ is a closed prequantization over $(W,\omega^{\prime})$, then for any $k\in\mathbb{N}_{>0}$ we denote by $M_{k}:=M_{1}/\mathbb{Z}_{k}$, where $\mathbb{Z}_{k}$ is seen as a subgroup of $\mathbb{T}^{1}$ via the map $[j]\mapsto\frac{2\pi j}{k}$, and $\xi^{\prime}_{k}$ the contact distribution on $M_{k}$ given by the kernel of the unique $1$-form $\alpha^{\prime}_{k}$ whose pull back via the covering $M_{1}\to M_{1}/\mathbb{Z}_{k}$ is $\alpha^{\prime}_{1}$. Finally we denote by $\pi^{\prime}_{k}:M_{k}\to W$ the natural projection. Note that ${\pi^{\prime}_{k}}^{*}\omega^{\prime}=\mathrm{d}\alpha^{\prime}_{k}$ and that $(M_{k},\xi^{\prime}_{k})$ is a prequantization over $(W,k\omega^{\prime})$ by taking $k\alpha^{\prime}_{k}$ to be the contact form inducing the Zoll Reeb flow of minimal period $2\pi$.
Theorem 1.3.
Suppose that $(M_{1},\xi^{\prime}_{1})$ is a strongly orderable prequantization over a closed symplectic manifold $(W,\omega^{\prime})$ and let $\Psi_{1}:(\mathbb{B}^{2n}(R_{1}),\omega_{0})\hookrightarrow(W,\omega^{\prime})$, $\Psi_{2}:(\mathbb{B}^{2n}(\lambda R_{2}),\omega_{0})\hookrightarrow(W,\omega^{\prime})$ be two symplectic embeddings for some $R_{1}\in(0,\sqrt{2}]$, $R_{2}>0$, $\lambda>\sqrt{3}$, and $a_{1}\in(0,R_{1}],\ a_{2}\in(0,R_{2}]$ be two positive real numbers. For any $k\in\mathbb{N}_{>0}$, if there exists $j\in\mathbb{N}_{>0}$ such that
$\pi a_{1}^{2}>\frac{2\pi}{k}\cdot j\geq\pi a_{2}^{2},$
then $B^{1}_{k}(a_{1})$ cannot be contactly squeezed inside $B^{2}_{k}(a_{2})$, where by definition $B^{i}_{k}(a_{i}):={\pi^{\prime}_{k}}^{-1}(\Psi_{i}(\mathbb{B}^{2n}(a_{i})))\subset(M_{k},\xi^{\prime}_{k})$ for all $i\in\{1,2\}$.
As for Theorem 1.1 we prove Theorem 1.3 by constructing a contact capacity defined on open subsets of $(M_{k},\xi^{\prime}_{k})$ and by estimating this contact capacity for the lift of Darboux balls. The contact capacity this time will come from the order spectral selectors defined in [4] on the universal cover of the isotopy class of the Legendrian $\Delta_{k}$ lying in the contact product of $(M_{k},\xi^{\prime}_{k})$ with itself.
Remark 1.4.
  1. (1)
    In Theorem 1.3 we make the assumption that $R_{1}\leq\sqrt{2}$ so that $\pi R_{1}^{2}$ is not greater than the minimal period of the Reeb flow associated to the contact form $\alpha^{\prime}_{1}$ satisfying ${\pi^{\prime}_{1}}^{*}\omega^{\prime}=\mathrm{d}\alpha^{\prime}_{1}$ (see the 4th step of the proof of Proposition 8.2 for more details). This technical assumption was not made in Theorem 1.1 because, by Theorem [27, Thm 1] discussed in Remark 1.2, it is automatically satisfied. In particular, when $(M_{1},\xi^{\prime}_{1})$ is a strongly orderable prequantization, one can expect to detect non-squeezing phenomena in $(M_{k},\xi^{\prime}_{k})$ thanks to Theorem 1.3 only for $k>1$.
  2. (2)
    Even if the lens space $(L_{k}^{2n+1},\xi_{k})$ defined in Section 1.1 can be seen as a prequantization over $(\mathbb{C}\mathrm{P}^{n},k\omega=2k\omega_{FS})$, one cannot deduce Theorem 1.1 from Theorem 1.3. First, it is not clear to the author whether contact lens spaces are strongly orderable or not in general. Indeed $(L_{k}^{2n+1},\xi_{k})$, for $n>1$, is not hypertight since it satisfies the Weinstein conjecture (see [22]), i.e. the Reeb flow of any contact form $\alpha$ satisfying $\ker\alpha=\xi_{k}$ admits a periodic orbit, and its fundamental group $\pi_{1}(L_{k}^{2n+1})$ is finite. Moreover, if $n\in\mathbb{N}_{>0}$ is not a power of $2$, Zhou [39] showed that $(L_{2}^{2n-1},\xi_{2})$ is not Liouville fillable. Second, for the reasons discussed in the first point of this Remark 1.4 and the fact that $(L_{1}^{2n+1},\xi_{1})=(\mathbb{S}^{2n+1},\xi_{1})$ is not strongly orderable for $n\geq 1$ (see [15]), Theorem 1.3 cannot detect non-squeezing phenomena in $(L_{p}^{2n+1},\xi_{p})$ when $p$ is a prime number even when it is known that the latter is strongly orderable. In particular, $(L_{2}^{3},\xi_{2})$ is strongly orderable, since the standard symplectic codisk bundle of the $2$-sphere is a Liouville filling whose symplectic homology does not vanish (see Example 1.5 below), but from what we have just said Theorem 1.3 cannot detect non-squeezing phenomena in the latter contact manifold while Theorem 1.1 does (see Remark 1.2).
  3. (3)
    We conjecture that Theorem 1.3 also holds if we assume $(M_{1},\xi^{\prime}_{1})$ to be orderable in the sense of [16], i.e. there does not exist any positive contractible loop of contactomorphisms, which is a priori weaker than to assume $(M_{1},\xi^{\prime}_{1})$ to be strongly orderable. The validity of this conjecture would be an easy consequence of the proof given below in Section 8 together with the validity of the conjecture made in [3] about the spectrality of the (pseudo-)spectral selectors for contactomorphisms.
  4. (4)
    It seems very plausible that with the tools developped in [6] one could derive similar non-squeezing phenomena satisfied by the ideal contact boundary of a semipositive and convex-at-infinity symplectic manifold $W$ whose unit in symplectic homology is not eternal.
Examples 1.5.
  1. (1)
    For any closed manifold $X$, its associated cosphere bundle with its standard contact structure is strongly orderable by [7, Thm. 1.18], since it is Liouville fillable by the standard symplectic codisk bundle and the symplectic homology of the latter does not vanish (see for instance the introduction of [7] or [1, Example 1.6] combined with [34, Thm. 13.3]). If we assume moreover that $X$ admits a Riemaniann metric $g$ so that all its geodesics are closed with length $2\pi$ (see [32, Chap. 3.4] for examples) then the cosphere bundle is, in addition to be strongly orderable, also a prequantization (see [20, Thm. 7.2.5]).
  2. (2)
    If $(W,\omega^{\prime})$ is a closed symplectic manifold so that $[\omega^{\prime}],H_{2}(W,\mathbb{Z})\right>\subset 2\pi\mathbb{Z}$ and $\pi^{\prime}_{1}:M_{1}\to W$ denotes the $\mathbb{R}/2\pi\mathbb{Z}$-principal bundle whose Euler class, after taking its image by the map $H^{2}(W,\mathbb{Z})\to H^{2}(W,\mathbb{R})\simeq H^{2}_{\mathrm{dR}}(W,\mathbb{R})$, is equal to $-[\frac{\omega^{\prime}}{2\pi}]$, then $M_{1}$ carries a contact form $\alpha^{\prime}_{1}$ that turns $(M_{1},\ker\alpha^{\prime}_{1})$ into a prequantization over $(W,\omega^{\prime})$, i.e. ${\pi_{1}^{\prime}}^{*}\omega^{\prime}=d\alpha_{1}^{\prime}$, the Reeb flow of $\alpha_{1}^{\prime}$ is Zoll of minimal period $2\pi$ and it induces the $\mathbb{R}/2\pi\mathbb{Z}$ action (see for instance [20, Chap 7.2]). If we suppose in addition that $(W,\omega^{\prime})$ is symplectically aspherical, i.e. $[\omega^{\prime}],H_{2}^{S}(M,\mathbb{Z})\right>=0$ where $H_{2}^{S}(M,\mathbb{Z})$ is the image of the Hurewitz homomorphism $\pi_{2}(M)\to H_{2}(M,\mathbb{Z})$, then $(M_{1},\xi^{\prime}_{1})$ is hypertight and in particular is strongly orderable thanks to [7, Thm. 1.19]. Therefore we get non-squeezing phenomena thanks to Theorem 1.3 in $(M_{k},\xi^{\prime}_{k})$ for $k>1$ big enough.
1.3. Discussion
It is interesting to note that the $\alpha_{k}$, and $\alpha^{\prime}_{k}$ Reeb flows on $(L_{k}^{2n+1},\xi_{k})$ and on $(M_{k},\xi^{\prime}_{k})$, as described above in Sections 1.1 and 1.2, are periodic with period $\frac{2\pi}{k}$ for all integers $k\geq 1$. Indeed, the periodicity of the Reeb flow often plays a crucial role in the existence of non trivial global invariants and this article is no exception to this fact (see Lemma 2.2). Furthermore, the global invariants one can deduce from Theorem 1.1 and Theorem 1.3 naturally take value in the discrete set consisting in multiples of this period. Similarly, in the contact manifold $\left(\mathbb{R}^{2n}\times\mathbb{R}/\mathbb{Z},\ker(\alpha_{0}:=\mathrm{d}z-\frac{1}{2}\sum\limits_{i=1}^{n}(x_{i}dy_{i}-y_{i}dx_{i}))\right)$ where the first contact non-squeezing phenomena was described by Eliashberg, Kim and Polterovich in [15], the $\alpha_{0}$-Reeb flow is $1$-periodic, and they show that it is impossible to contactly squeeze $\mathbb{B}^{2n}(A)\times\mathbb{R}/\mathbb{Z}$ inside $\mathbb{B}^{2n}(a)\times\mathbb{R}/\mathbb{Z}$ if there exists an integer $j\in\mathbb{N}_{>0}$ so that $A>j\geq a$.
The same kind of phenomenon can be observed when studying bi-invariant metrics on the group $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ or its universal cover $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(M,\xi)$. While Sandon in [35] constructs an unbounded bi-invariant metric on $\mathrm{Cont}_{0}^{\mathrm{c}}(\mathbb{R}^{2n}\times\mathbb{R}/\mathbb{Z},\ker\alpha_{0})$ taking integer values, the authors in [4] construct an unbounded bi-invariant metric on $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ taking values in $\frac{2\pi}{k}\mathbb{N}$ [^4] It is shown in [18] that, for any contact manifold $(M,\xi)$, a bi-invariant metric on $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ (resp. $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(M,\xi)$) has to induce the discrete topology on $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ (resp. on $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(M,\xi)\setminus\pi_{1}(\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi))$).
What we have just described should be nuanced: beeing a contact manifold admitting a periodic Reeb flow is not a sufficient condition to have rigidity results and existence of global invariants. In fact, the $\alpha_{1}$-Reeb flow of the standard contact sphere $(\mathbb{S}^{2n+1},\ker\alpha_{1})$ is periodic but any two non-empty open domains that are not the whole sphere can be contactly squeezed inside each other. In the same way, it is shown in [18] that, for $n\in\mathbb{N}_{>0}$, the groups $\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\ker\alpha_{1})$ and $\widetilde{\mathrm{Cont}_{0}}(\mathbb{S}^{2n+1},\ker\alpha_{1})$ are bounded, i.e. any bi-invariant metric on these groups has to be bounded.
It seems that another important property that a contact manifold should have so that one could expect meaningful invariant measurements is to be orderable (see 2.2, 7.1 for definitions). For $n>1$ and $k\geq 2$, the sphere $(\mathbb{S}^{2n+1},\ker\alpha_{1})$ and the lens space $(L_{k}^{2n+1},\ker\alpha_{k})$ are perfect representative cases since the sphere is not orderable [15] while the lens space is [22], and, as we have just mentionned, interesting contact measurements exist for the latter but not for the former. In the same direction, a similar observation can be made for bi-invariant metrics on contactomorphisms groups. On the one hand, to the knowledge of the author, the only examples in the litterature of closed contact manifolds $(M,\xi)$ admitting unbounded bi-invariant metrics on $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ are the contact manifolds $(M,\xi)$ that are orderable and that admit a periodic Reeb flow [18, 11, 22, 4]. On the other hand, for a large family of closed non-orderable contact manifolds $(M,\xi)$ Courte and Massot show that $\mathrm{Cont}_{0}(M,\xi)$ and $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ have to be bounded [12].
Actually, while convenient but not sufficient, the existence of a periodic Reeb flow is not even a necessary condition to have contact global invariants. For example, the unitary cotangent of the torus $\mathbb{P}^{+}T^{*}\mathbb{T}^{n}$ endowed with its canonical contact structure $\xi_{\mathrm{can}}$ does not admit any periodic Reeb flow when $n>1$ (see for example [8]). But, as briefly mentionned above, one can use an interesting contact global invariant for domains in $(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$, which is called the shape555unlike the other invariant measurements discussed so far, the shape can be used to construct “non-discrete” measurements, to deduce non-squeezing phenomena and other rigidity results concerning this contact manifold [14, 5, 33, 16]. It would therefore be interesting to investigate whether the groups $\widetilde{\mathrm{Cont}_{0}}(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$ and $\ \mathrm{Cont}_{0}(\mathbb{P}^{+}T^{*}\mathbb{T}^{n},\xi_{\mathrm{can}})$ are bounded or not (see also [3, Prop. 4.14, Rem. 4.15]).
Organization of the paper
In Section 2 we set some conventions and give basic definitions and properties of orderability and spectral selectors. In Section 3 we briefly recall the construction and main properties of the spectral selectors on $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ constructed in [4]. In Section 4 we define the above mentionned contact capacity for open domains of $(L_{k}^{2n+1},\xi_{k})$ coming from the spectral selectors. In this same Section 4 we state - without prooving - some properties of this contact capacity (see Theorem 4.1) which allow us to give the proof of Theorem 1.1. In Section 5 we show a capacity-energy inequality satisfied by the contact capacity. In Section 6 we prove Theorem 4.1 by estimating and computing the contact capacity of domains that are lifts of Darboux balls to $(L_{k}^{2n+1},\xi_{k})$ using the capacity-energy inequality of Section 5 and some explicit constructions. In Section 7 we define the spectral selectors on $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ when $(M,\xi)$ is a strongly orderable closed contact manifold. Finally in Section 8 we use these spectral selectors to define a contact capacity for domains in strongly orderable prequantizations and prove Theorem 1.3.
Acknowledgment
The author thanks Alberto Abbondandolo, Johanna Bimmermann, Stefan Nemirovski and Sheila Sandon for interesting and stimulating discussions.
The author was partially supported by the Deutsche Forschungsgemeinschaft under the Collaborative Research Center SFB/TRR 191 - 281071066 (Symplectic Structures in Geometry, Algebra and Dynamics).
2. Spectral Selectors
2.1. Conventions and definition
Let $(M,\xi)$ be a closed coorientable contact manifold. We say that a $[0,1]$-family of contactomorphisms $\{\phi_{t}\}_{t\in[0,1]}\subset\mathrm{Cont}(M,\xi)$, that we abbreviate with $(\phi_{t})$, is a contact isotopy if the map $[0,1]\times M\to M$, $x\mapsto\phi_{t}(x)$ is smooth. We denote by $\mathrm{Cont}_{0}(M,\xi)$ the group of contactomorphisms $\phi$ that are contact isotopic to the identity: there exists a contact isotopy $(\phi_{t})$ such that $\phi_{0}=\mathrm{id}$ and $\phi_{1}=\phi$. We endow $\mathrm{Cont}_{0}(M,\xi)$ with the $C^{1}$-topology, denote by $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ its universal cover and $\Pi:\widetilde{\mathrm{Cont}_{0}}(M,\xi)\to\mathrm{Cont}_{0}(M,\xi)$ the projection. We often see $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ as equivalence classes of contact isotopies starting at the identity where two isotopies represent the same equivalence class if they are homotopic relatively to endpoints.
A point $x\in M$ is said to be a discriminant of a contactomorphism $\phi\in\mathrm{Cont}_{0}(M,\xi)$ if $\phi(x)=x$ and $(\phi^{*}\alpha)_{x}=\alpha_{x}$, for some - and thus all - contact form $\alpha$ supporting $\xi$, i.e. $\ker\alpha=\xi$.
Let $\alpha$ be a contact form supporting $\xi$, i.e. $\ker\alpha=\xi$. We say that a point $x\in M$ is an $\alpha$-translated point of $\phi\in\mathrm{Cont}_{0}(M,\xi)$ if there exists $T\in\mathbb{R}$ so that $x$ is a discriminant point of $\phi_{\alpha}^{-T}\circ\phi$, where $\phi_{\alpha}^{s}$ denotes the Reeb flow at time $s\in\mathbb{R}$ of $\alpha$, i.e. the flow of the vector field $R_{\alpha}$ uniquely defined by the two equations $\alpha(R_{\alpha})\equiv 1$ and $\mathrm{d}\alpha(R_{\alpha},\cdot)\equiv 0$. For any $\phi\in\mathrm{Cont}_{0}(M,\xi)$ we denote and define its $\alpha$-spectrum by
$\mathrm{Spec}^{\alpha}(\phi):=\left\{t\in\mathbb{R}\ |\ \phi_{\alpha}^{-t}\circ\phi\text{ has a discriminant point}\right\}.$
Similarly, the $\alpha$-spectrum $\mathrm{Spec}^{\alpha}(\widetilde{\varphi})$ of any $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ is by definition
$\mathrm{Spec}^{\alpha}(\widetilde{\varphi}):=\mathrm{Spec}^{\alpha}\left(\Pi(\widetilde{\varphi})\right).$
Definition 2.1.
We say that a map $c:\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)\to\mathbb{R}$ is a $\alpha$-spectral selector if it is $C^{1}$-continuous and if $c(\widetilde{\varphi})\in\mathrm{Spec}^{\alpha}(\widetilde{\varphi})$.
2.2. Orderability, Periodic Reeb flow and invariance of spectral selectors
Let $(M,\xi)$ be a closed coorientable contact manifold. From now on we fix a coorientation of $\xi$, i.e. a vector field $X\in\chi(M)$ so that $X(M)\subset TM\setminus\xi$. We say that a contact form $\alpha$ supports $\xi$ if $\ker\alpha=\xi$ and $\alpha(X)>0$.
A contact isotopy $(\phi_{t})$ is said to be non-negative (resp. positive) if for some, and thus all, contact form $\alpha$ supporting $\xi$ we have $\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\geq 0$ (resp. $>0$) for all $x\in M$. Following [16] we define a binary relation on $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ by saying that $\widetilde{\varphi}\preceq\widetilde{\psi}$ if there exists a path $(\widetilde{\varphi}_{t})$ joining them such that $(\Pi(\widetilde{\varphi}_{t}))$ is a non-negative contact isotopy. It is shown in [16] that this binary relation is a partial order if and only there does not exist any positive contractible loop of contactomorphisms, and in this case we say that $(M,\xi)$ is orderable. The following Lemma will be crucial in this paper and was, to some extent, already noticed by Sandon in [36]
Lemma 2.2.
Let $(M,\xi)$ be a orderable closed contact manifold and $\alpha$ be a contact form supporting $\xi$ whose Reeb flow is periodic of minimal period $T>0$. If $c:\widetilde{\mathrm{Cont}_{0}}(M,\xi)\to\mathbb{R}$ is a $\alpha$-spectral selector which is monotone, i.e. $\widetilde{\varphi}_{1}\preceq\widetilde{\varphi}_{2}$ implies that $c(\widetilde{\varphi}_{1})\leq c(\widetilde{\varphi}_{2})$, then
$\lceil c(\widetilde{\psi}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}^{-1})\rceil_{T}=\lceil c(\widetilde{\varphi})\rceil_{T}\quad\quad\text{ for all }\widetilde{\varphi},\ \widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi).$
Proof.
Consider $(\psi_{t})$ a path representing $\widetilde{\psi}$ and denote by $\widetilde{\psi}_{s}:=[(\psi_{st})]$ for all $s\in[0,1]$. Let us first suppose that $\Pi(\widetilde{\varphi})$ does not have discriminant point. Thus $\Pi(\widetilde{\psi}_{s}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{s}^{-1})$ does not have discriminant point and so $c(\widetilde{\psi}_{s}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{s}^{-1})\notin T\mathbb{Z}$ for all $s\in[0,1]$. We then conclude using the continuity of $s\mapsto c(\widetilde{\psi}_{s}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}_{s}^{-1})$. We treat the general case using the fact that $\mathrm{Spec}^{\alpha}(\widetilde{\varphi})$ is nowhere dense (see [1, Lemma 2.11]) and therefore there exists a sequence $(\varepsilon_{n})_{n\in\mathbb{N}}$ of positive real numbers going to 00 so that the sequence $\Pi(\widetilde{\phi_{\alpha}^{-\varepsilon_{n}}}\cdot\widetilde{\varphi})$ does not have any discriminant for all $n\in\mathbb{N}$. Since $\widetilde{\phi_{\alpha}^{-\varepsilon_{n}}}\cdot\widetilde{\varphi}\preceq\widetilde{\varphi}$ and the sequence $(\widetilde{\phi_{\alpha}^{-\varepsilon_{n}}}\cdot\widetilde{\varphi})$ converges to $\widetilde{\varphi}$ in the $C^{1}$-topology we conclude using the previous case and the monotonicity of the $\alpha$-selector. ∎
3. Spectral selectors for lens spaces
In [4] the authors construct a $\alpha_{k}$-spectral selector $c_{k}:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\ker\alpha_{k})\to\mathbb{R}$ for any $n\in\mathbb{N}_{>0}$ and $k\geq 2$ (denoted in [4] by $c_{0}$ or $c_{+}$). Its construction, that we present briefly below, is based on generating functions techniques and Givental’s non-linear Maslov index [22, 21]. From now on we fix $n\in\mathbb{N}_{>0}$ and an integer $k\geq 2$.
3.0.1. Givental’s non-linear Maslov index
The Givental’s non-linear Maslov index $\mu:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{Z}$ is a quasimorphism that detects discriminant points in the following sense: if $\{\widetilde{\varphi}_{t}\}_{t\in[0,1]}\subset\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ is a smooth path, i.e. $(\phi_{t}:=\Pi(\widetilde{\varphi}_{t}))$ is a contact isotopy, such that $\mu(\widetilde{\varphi}_{0})\neq\mu(\widetilde{\varphi}_{1})$, then there exists $t_{0}\in[0,1]$ such that $\Pi(\widetilde{\varphi}_{t_{0}})$ has a discriminant point. Denoting by $\widetilde{\phi_{\alpha_{k}}^{T}}:=[(\phi_{\alpha_{k}}^{tT})]\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$, for $T\in\mathbb{R}$, it is shown moreover in [21, 22, 4] that $\mu$ has the following properties.
Lemma 3.1 ([21, 22, 4]).
For all $\widetilde{\varphi},\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
  1. (1)
    the map $\mathbb{R}\to\mathbb{Z}$, $T\mapsto\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})$ is lower semi-continuous
  2. (2)
    if $\mu(\widetilde{\varphi})$ is even then $\mu(\widetilde{\psi}\cdot\widetilde{\varphi})\leq\mu(\widetilde{\psi})+\mu(\widetilde{\varphi})$
  3. (3)
    $\mu(\widetilde{\phi_{\alpha_{k}}^{T}})=2(n+1)\left\lceil\frac{T}{2\pi}\right\rceil$ for all $T\in\mathbb{R}$
  4. (4)
    $\mu(\widetilde{\varphi})\leq\mu(\widetilde{\psi})$ if $\widetilde{\varphi}\preceq\widetilde{\psi}$,
where $\preceq$ is the bi-invariant binary relation introduced by Eliashberg-Polterovich in [16], i.e. $\widetilde{\varphi}\preceq\widetilde{\psi}$ if $\widetilde{\psi}\cdot\widetilde{\varphi}^{-1}$ can be represented by a contact isotopy $(\phi_{t})$ such that $\alpha_{k}\left(\frac{d}{dt}\phi_{t}\right)\geq 0$.
Let us give a brief overview of its construction. To do consider $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ and $(\phi_{t})$ any contact isotopy representing it.
Firstly, let us denote by $(\overline{\phi_{t}})\subset\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\xi_{1})$ the lifted $\mathbb{Z}_{k}$-equivariant contact isotopy starting at the identity associated to the contact isotopy $(\phi_{t})$, i.e. $\pi_{1}^{k}\circ\overline{\phi_{t}}=\phi_{t}\circ\pi_{1}^{k}$ for all $t\in[0,1]$ and $\overline{\phi_{0}}=\mathrm{id}$ where $\pi_{1}^{k}:\mathbb{S}^{2n+1}\to L_{k}^{2n+1}$ is the covering map. We want to construct a path of $C^{1}$-functions $(f_{t}:L_{k}^{2(n+N)+1}\to\mathbb{R})$, for some integer $N\geq 0$, so that for all $t\in[0,1]$
$\{x\in L_{k}^{2(n+N)+1}\ |\ \mathrm{d}_{x}f_{t}=0,\ f_{t}(x)=0\}\overset{1-1}{\leftrightarrow}\{\pi_{1}^{k}(x)\in L_{k}^{2n+1}\ |\ x\text{ is a disc. point of }\overline{\phi}_{t}\},$
i.e. there is a one-to-one correspondance between the critical points of $f_{t}$ with critical value 0 and the discriminant points of $\phi_{t}$ that lift to a $\mathbb{Z}_{k}$-family of discriminant points of $\overline{\phi}_{t}$. This path of function $(f_{t})$ is constructed in the following way. Identifying the symplectization of $(\mathbb{S}^{2n+1},\xi_{1})$ with $(\mathbb{R}^{2n+2}\setminus\{0\},\omega_{0})$, the $\mathbb{Z}_{k}$-equivariant contact isotopy $(\overline{\phi}_{t})\subset\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\xi_{1})$, can itself be uniquely lifted to a $\mathbb{Z}_{k}\times\mathbb{R}_{>0}$-equivariant Hamiltonian isotopy $(\psi_{t})\subset\mathrm{Ham}(\mathbb{R}^{2n+2}\setminus\{0\},\omega_{0})$
$\psi_{t}(z)=e^{-\frac{1}{2}g_{t}(z/\|z\|)}\|z\|\overline{\phi}_{t}\left(\frac{z}{\|z\|}\right),\quad\text{ for all }t\in[0,1],\ z\in\mathbb{R}^{2n+2}\setminus\{0\},$
where $g_{t}:\mathbb{S}^{2n+1}\to\mathbb{R}$ is the $\alpha_{1}$-conformal factor of $\overline{\phi}_{t}$, i.e. $\overline{\phi}_{t}^{*}\alpha_{1}=e^{g_{t}}\alpha_{1}$. Extending $\psi_{t}$ to the whole $\mathbb{R}^{2n+2}$ by putting $\psi_{t}(0)=0$ the authors in [22, Prop 2.2 and Prop 2.14] show the existence of an integer $m\in\mathbb{N}$ and a path of $C^{1}$-functions $(F_{t}:\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}\to\mathbb{R})$, where $2N=(2n+2)\cdot 2m$, such that denoting by $\Sigma_{F_{t}}:=\{(x,v)\in\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}\ |\ \partial_{v}F(x,v)\equiv 0\in T^{*}\mathbb{R}^{2n}\}$ the fiber critical points of $F_{t}$ we have666 in the following $\mathbb{R}^{2j}$ is identified with $\mathbb{C}^{j}$, for all $j\in\mathbb{N}_{>0}$, as in the introduction and $T^{*}\mathbb{R}^{2n+2}$ with $T\mathbb{R}^{2n+2}\simeq\mathbb{R}^{2n+2}\times\mathbb{R}^{2n+2}$ using the standard Euclidean metric
  1. (1)
    $i_{F_{t}}:\Sigma_{F_{t}}\to T^{*}\mathbb{R}^{2n+2},\ (x,v)\hookrightarrow\left(x,\partial_{x}F_{t}(x,v)\right)$ is injective and moreover
    $i_{F_{t}}(\Sigma_{F_{t}})=\tau(\mathrm{graph}(\psi_{t}))$
    where $\text{graph}(\psi_{t}):=\{(x,\psi_{t}(x))\ |\ x\in\mathbb{R}^{2n+2}\}$ and $\tau(z,Z)=\left(\frac{z+Z}{2},i(z-Z)\right)$ for all $z,Z\in\mathbb{R}^{2n+2}\times\mathbb{R}^{2n+2}.$
  2. (2)
    $F_{t}(\lambda e^{\frac{2i\pi j}{k}}z)=\lambda^{2}F_{t}(z)$ for all $\lambda>0$, $z\in\mathbb{R}^{2n+2}\times\mathbb{R}^{2N}$ and $j\in\mathbb{N}$
  3. (3)
    $F_{0}:\mathbb{R}^{2n+2}\times(\mathbb{R}^{2n+2})^{2m}\to\mathbb{R}$, $(z,(Z_{1},\cdots,Z_{2m}))\mapsto-2\sum\limits_{j=1}^{m}\left<Z_{2j}-z,i(Z_{2j-1}-z)\right>$.
The path of functions $(F_{t})$ induces the desired path $(f_{t}:L_{k}^{2(n+N)+1}\to\mathbb{R})$ and $(F_{t})$ is called a based family of generating functions associated to $(\phi_{t})$.
Secondly, suppose that $k=p\in\mathbb{N}_{>0}$ is a prime number. If $(\phi^{\prime}_{t})$ is another contact isotopy in $\mathrm{Cont}_{0}(L_{p}^{2n+1},\xi_{p})$ representing the same element $\widetilde{\varphi}$ and if $(f_{t}:L_{p}^{2N+1}\to\mathbb{R})$ and $(f_{t}^{\prime}:L_{p}^{2N^{\prime}+1}\to\mathbb{R})$ are two paths of functions that are induced by two based families of generating functions associated respectively to $(\phi_{t})$ and $(\phi^{\prime}_{t})$, then it is proven in [22, Prop. 2.20 and Section 4] that
$\operatorname{ind}(\{f_{0}\leq 0\})-\operatorname{ind}(\{f_{1}\leq 0\})=\operatorname{ind}(\{f^{\prime}_{0}\leq 0\})-\operatorname{ind}(\{f_{1}^{\prime}\leq 0\}).$
Here the index $\operatorname{ind}(A)$ of an open subset $A\subset L_{p}^{j}$, for any integer $j\geq 3$, is the dimension over the field $\mathbb{Z}_{p}$ of the image of $i_{A}^{*}:H^{*}(L_{p}^{j},\mathbb{Z}_{p})\to H^{*}(A,\mathbb{Z}_{p})$, where $i_{A}$ denotes the inclusion and $i_{A}^{*}$ the induced map on singular homology, and the index of a closed subset $F$ is by definition $\operatorname{ind}(F):=\min\{\operatorname{ind}(A)\ |\ A\text{ open and }F\subset A\}$. The Givental’s non-linear Maslov index of $\widetilde{\varphi}$ is then defined to be
$\mu(\widetilde{\varphi}):=\operatorname{ind}(\{f_{0}\leq 0\})-\operatorname{ind}(\{f_{1}\leq 0\}).$
Finally, for a general integer $k\geq 2$, denoting by $p\in\mathbb{N}_{>0}$ the smallest prime dividing $k$, we define $\mu(\widetilde{\varphi})$ by
$\mu(\widetilde{\varphi}):=\mu(\mathcal{L}_{k}^{p}(\widetilde{\varphi})),$
where $\mathcal{L}_{k}^{p}:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\widetilde{\mathrm{Cont}_{0}}(L_{p}^{2n+1},\xi_{p})$ denotes the natural lift map.
3.0.2. The spectral selector on $(L_{k}^{2n+1},\xi_{k})$
The $\alpha_{k}$-spectral introduced in [4] is defined by
$c_{k}:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{R}\quad\quad\widetilde{\varphi}\mapsto\inf\left\{T\in\mathbb{R}\ |\ \mu\left(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi}\right)\leq 0\right\}.$
Before stating its main properties we introduce some notations. For any $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ we consider the following set
$\overline{\mathrm{Spec}^{\alpha_{k}}}(\widetilde{\varphi}):=\mathrm{Spec}^{\alpha_{1}}(\overline{\widetilde{\varphi}})\subset\mathrm{Spec}^{\alpha_{k}}(\widetilde{\varphi})$
where $\overline{\widetilde{\varphi}}\in\widetilde{\mathrm{Cont}_{0}}(\mathbb{S}^{2n+1},\xi_{1})$ denotes the $\mathbb{Z}_{k}$-equivariant lift of $\widetilde{\varphi}$. For any positive number $T>0$ and any $x\in\mathbb{R}$ we denote by $\lceil x\rceil_{T}:=T\left\lceil\frac{x}{T}\right\rceil$ the smallest multiple of $T$ greater or equal to $x$.
Proposition 3.2 ([4]).
The map $c_{k}:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{R}$ is a $\alpha_{k}$-spectral selector that moreover satisfies for all $\widetilde{\varphi},\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
  1. (1)
    $c_{k}(\mathrm{id})=0$
  2. (2)
    $\widetilde{\varphi}\preceq\widetilde{\psi}$ implies that $c_{k}(\widetilde{\varphi})\leq c_{k}(\widetilde{\psi})$
  3. (3)
    $\mathrm{id}\preceq\widetilde{\varphi}$ and $\widetilde{\varphi}\neq\mathrm{id}$ implies that $c_{k}(\widetilde{\varphi})>0$
  4. (4)
    $c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi})\leq c_{k}(\widetilde{\psi})+\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$
  5. (5)
    for any contact isotopy $(\phi_{t})$ representing $\widetilde{\varphi}$
    $\int_{0}^{1}\min\alpha_{k}\left(\frac{d}{dt}\phi_{t}\right)\ \mathrm{d}t\leq c_{k}(\widetilde{\varphi})\leq\int_{0}^{1}\max\alpha_{k}\left(\frac{d}{dt}\phi_{t}\right)\ \mathrm{d}t$
  6. (6)
    $c_{k}(\widetilde{\varphi})\in\overline{\mathrm{Spec}^{\alpha_{k}}}(\widetilde{\varphi})$
  7. (7)
    $\lceil c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi}\cdot\widetilde{\psi}^{-1})\rceil_{\frac{2\pi}{k}}=\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$.
Before concluding this section it is worth noticing that we have a stronger triangular inequality and invariance by conjugation for elements in
[…]
Proposition 3.3.
For any $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(L_{k}^{2n+1},\xi_{k})$, $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
$c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi})\leq c_{k}(\widetilde{\psi})+c_{k}(\widetilde{\varphi})$
To prove Proposition 3.3, let us recall that $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is said to be $\alpha_{k}$-nondegenerate if for all $\alpha_{k}$-translated points $x\in L_{k}^{2n+1}$ of $\phi$ there does not exist any $X\in T_{x}L_{k}^{2n+1}\setminus\{0\}$ and $T\in\mathbb{R}$ such that
$\mathrm{d}_{x}(\phi_{\alpha_{k}}^{-T}\circ\phi)(X)=X\text{ and }\mathrm{d}_{x}g(X)=0$
where $g:L_{k}^{2n+1}\to\mathbb{R}$ is the $\alpha_{k}$-conformal factor of $\phi$, i.e. $\phi^{*}\alpha_{k}=e^{g}\alpha_{k}$. The next Lemma is also contained in [4, Thm 1.1].
Lemma 3.4.
Let $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$. If $\Pi(\widetilde{\varphi})$ is $\alpha_{k}$-nondegenerate then $\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})=0$ where $T:=c_{k}(\widetilde{\varphi})$.
Proof of Proposition.
3.3 Let us first suppose that $\Pi(\widetilde{\varphi})$ is $\alpha_{k}$-nondegenerate and denote by $T:=c_{k}(\widetilde{\varphi})$, $T^{\prime}:=c_{k}(\widetilde{\psi})$. Then by Lemma 3.4 we have that $\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})=0$ is even. Therefore, since $\widetilde{\psi}$ commutes with $\widetilde{\phi_{\alpha_{k}}^{s}}$ for all $s\in\mathbb{R}$ we have
$\mu(\widetilde{\phi_{\alpha_{k}}^{-(T^{\prime}+T)}}\cdot\widetilde{\psi}\cdot\widetilde{\varphi})=\mu(\widetilde{\phi_{\alpha_{k}}^{-T^{\prime}}}\cdot\widetilde{\psi}\cdot\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})$ $\leq\mu(\widetilde{\phi_{\alpha_{k}}^{-T^{\prime}}}\cdot\widetilde{\psi})+\mu(\widetilde{\phi_{\alpha_{k}}^{-T}}\cdot\widetilde{\varphi})$
$\leq\mu(\widetilde{\phi_{\alpha_{k}}^{-T^{\prime}}}\cdot\widetilde{\varphi})\leq 0,$
where the first inequality follows from the 2nd point of Lemma 3.1 and the last inequality from the 1st property of Lemma 3.1. This implies that
$c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi})\leq c_{k}(\widetilde{\psi})+c_{k}(\widetilde{\varphi})$
which finishes the proof in the case where $\Pi(\widetilde{\varphi})$ is non degenerate.
The general case can now be deduced by the $C^{1}$-continuity of $c_{k}$ and the $C^{1}$-density, inside $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$, of the set of elements $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ such that $\Pi(\widetilde{\psi})$ is $\alpha_{k}$-nondegenerate (it is a direct consequence of [4, Lemma 3.1] and the fact that $\Pi:\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is a local homeomorphism when endowing both spaces with the $C^{1}$-topology).∎
Remark 3.5.
A $\mathbb{T}^{1}$-equivariant contactomorphism $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ is always $\alpha_{k}$-degenerate since its $\alpha_{k}$-conformal factor is constantly equal to 00, i.e. $\phi^{*}\alpha_{k}=\alpha_{k}$, and $(\phi_{\alpha_{k}}^{-T}\circ\phi)^{*}R_{\alpha_{k}}=R_{\alpha_{k}}$ for all $T\in\mathbb{R}$.
4. Contact capacity in Lens spaces
For any non-empty open set $U\subset L_{k}^{2n+1}$ we denote by $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})$ the subgroup of $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ whose elements can be represented by contact isotopies $(\phi_{t})$ supported in $U$, i.e. $\overline{\underset{t\in[0,1]}{\bigcup}\mathrm{Supp}(\phi_{t})}\subset U$ where $\mathrm{Supp}(\phi_{t}):=\overline{\{x\in M\ |\ \phi_{t}(x)\neq x\}}$. Let $C_{k}:\{\text{non-empty open sets of }L_{k}^{2n+1}\}\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ be the map:
$U\mapsto C_{k}(U):=\sup\{c_{k}(\widetilde{\varphi})\ |\ \widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})\}.$
Theorem 4.1.
For all non-empty open sets $U,V\subset L_{k}^{2n+1}$ and $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$
  1. (1)
    (monotonicity) $C_{k}(U)\leq C_{k}(V)$ if $U\subset V$
  2. (2)
    (invariance by conjugation) $\lceil C_{k}(U)\rceil_{\frac{2\pi}{k}}=\lceil C_{k}(\phi(U))\rceil_{\frac{2\pi}{k}}$
  1. (3)
    (lower bound) if $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ is a symplectic embedding for some $R>0$ then $C_{k}(B_{k}(r))\geq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$,
  2. (4)
    (upper bound) if $\Psi:(\mathbb{B}^{2n}(\lambda R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ is a symplectic embedding for some $R>0$ and $\lambda>\sqrt{3}$ then $C_{k}(B_{k}(r))\leq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
The monotonicity property of $C_{k}$ in Theorem 4.1 follows by definition. The invariance by conjugation of $C_{k}$ follows from the invariance by conjugation of the map $\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})\to\mathbb{R},\ \widetilde{\varphi}\mapsto\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}$ (see Lemma 2.2). To give the complete proof of Theorem 4.1 it thus remains to prove the last two estimates. Let us admit them for the moment and deduce the proof of Theorem 1.1:
Proof of Theorem 1.1.
Suppose by contradiction that there exists $\phi\in\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ such that $\phi(B^{1}_{k}(a_{1}))\subset B^{2}_{k}(a_{2})$. Then, using the 1st and 4th point of Theorem 4.1, we have on one side
$C_{k}(\phi(B^{1}_{k}(a_{1})))\leq C_{k}(B^{2}_{k}(a_{2}))\leq\pi a_{2}^{2}\leq\frac{2\pi}{k}j.$ (1)
But on the other side using the 2nd and the 3rd point we get that
$\lceil C_{k}(\phi(B^{1}_{k}(a_{1})))\rceil_{\frac{2\pi}{k}}=\lceil C_{k}(B^{1}_{k}(a_{1}))\rceil_{\frac{2\pi}{k}}\geq\left\lceil\pi a_{1}^{2}\right\rceil_{\frac{2\pi}{k}}>\frac{2\pi}{k}j$
which contradicts (1) and finishes the proof of Theorem 1.1. ∎
We give the proof of the last two estimates of Theorem 4.1 in Section 6. But before, in the next Section 5 we prove a capacity-energy inequality satisfied by $C_{k}$ which will be crucial to prove the lower bound stated in Theorem 4.1.
5. Capacity-energy Inequality
Let $(M,\ker\alpha)$ be a closed contact manifold. For any subset $U\subset M$ we denote by $\mathcal{O}_{\alpha}(U):=\underset{t\in\mathbb{R}}{\bigcup}\phi_{\alpha}^{t}(U)$ its $\alpha$-orbit.
Definition 5.1.
A diffeomorphism $\phi$ of $M$ $\alpha$-displaces a subset $U\subset M$ if $\phi$ displaces the $\alpha$-orbit of $U$, i.e. $\phi(\mathcal{O}_{\alpha}(U))\bigcap\mathcal{O}_{\alpha}(U)=\emptyset.$
Remark 5.2.
  1. (1)
    If $(M,\xi)$ is a closed cooriented contact manifold, the Reeb flow of a generic contact form supporting $\xi$ has a dense orbit [10, Theorem C.4]. Hence, for such contact form $\alpha$, a diffeomorphism $\phi$ cannot displace the closure of the $\alpha$-orbit of any non-empty open set $U$, i.e. $\phi(\overline{\mathcal{O}_{\alpha}(U)})\cap\overline{\mathcal{O}_{\alpha}(U)}\neq\emptyset$.
  2. (2)
[…]
  3. (3)
    The first two remarks has to be put in contrast with the following fact: it has been shown in [37] that any contact manifold $(M,\xi)$ admits a contact form $\alpha$ and a non-empty open set $U\subset M$ where the $\alpha$-Reeb flow is $1$-periodic on $U$.
The aim of this Section is to prove the following:
Theorem 5.3.
Let $U\subset L_{k}^{2n+1}$ and $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$. If $\Pi(\widetilde{\psi})$ $\alpha_{k}$-displaces $U$ then $\lceil c_{k}(U)\rceil_{\frac{2\pi}{k}}\leq c_{k}(\widetilde{\psi}^{-1})+\lceil c_{k}(\widetilde{\psi})\rceil_{\frac{2\pi}{k}}$. If moreover $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(L_{k}^{2n+1},\xi_{k})$ then $c_{k}(U)\leq c_{k}(\widetilde{\psi}^{-1})+c_{k}(\widetilde{\psi}).$
Remark 5.4.
One can use [4], (see for example the proof of [4, Prop 5.3] combined with the Poincaré duality of [4, Thm 1.1]), to show moreover that for every $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$
$0\leq\lceil c_{k}(\widetilde{\psi})\rceil_{\frac{2\pi}{k}}+\lfloor c_{k}(\widetilde{\psi}^{-1})\rfloor_{\frac{2\pi}{k}}\leq 2\pi+\frac{2\pi}{k}.$
where $\lfloor x\rfloor_{\frac{2\pi}{k}}:=\frac{2\pi}{k}\left\lfloor x\cdot\frac{k}{2\pi}\right\rfloor$ is the biggest multiple of $\frac{2\pi}{k}$ smaller or equal to $x$ for any $x\in\mathbb{R}$. In particular if $U\subset L_{k}^{2n+1}$ is $\alpha_{k}$-displaceable then $\lceil c_{k}(U)\rceil_{\frac{2\pi}{k}}\leq 2\pi+\frac{4\pi}{k}$.
Theorem 5.3 will be a consequence of the next Proposition that can be stated for a general closed contact manifold $(M,\ker\alpha)$ admitting a $\alpha$-spectral selector. For any non-empty open set $U\subset M$, as before, $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi)\subset\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ denotes the subgroup of elements that can be represented by contact isotopies supported in $U$.
Proposition 5.5.
Let $U\subset M$ be a non-empty open set of a closed contact manifold $(M,\ker\alpha)$ and $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}^{c}(U,\xi)$. If $c:\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)\to\mathbb{R}$ is an $\alpha$-spectral selector and if $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ is such that $\Pi(\widetilde{\psi})$ $\alpha$-displaces $U\subset M$ then
$c(\widetilde{\psi}\cdot\widetilde{\varphi})=c(\widetilde{\psi}).$
Proposition is a consequence of the following:
Lemma 5.6.
Let $\widetilde{\psi},\widetilde{\varphi}\in\mathrm{Cont}_{0}(M,\xi)$. If $\Pi(\widetilde{\psi})$ $\alpha$-displaces the support of $\Pi(\widetilde{\varphi})$ then
$\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})\subset\mathrm{Spec}^{\alpha}(\widetilde{\psi}).$
Proof.
Let $t\in\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})$ and $x\in M$ a discriminant point of $\phi_{\alpha}^{-t}\circ(\psi\circ\phi)$ where $\psi:=\Pi(\widetilde{\psi})$ and $\phi:=\Pi(\widetilde{\varphi})$. In particular $\psi(\phi(x))=\phi_{\alpha}^{t}(x)$. Suppose by contradiction that $x$ lies in the support of $\phi$. Then $\phi(x)\in\mathrm{Supp}(\phi)\subset\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))$ and $\phi_{\alpha}^{t}(x)\in\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))$. But this contradicts the fact that $\psi(\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))\cap\mathcal{O}_{\alpha}(\mathrm{Supp}(\phi))=\emptyset$. Therefore $x$ is not in the support of $\phi$. Thus $x$ is a discriminant point of $\phi_{\alpha}^{-t}\circ\psi$ and $t\in\mathrm{Spec}^{\alpha}(\widetilde{\psi})$. In particular $\mathrm{Spec}^{\alpha}(\widetilde{\psi}\cdot\widetilde{\varphi})\subset\mathrm{Spec}^{\alpha}(\widetilde{\psi})$. ∎
Proof of Proposition 5.5.
Let $(\varphi_{t})_{t\in[0,1]}$ be a contact isotopy supported in $U$ representing $\widetilde{\varphi}$ and $\widetilde{\varphi}_{s}:=[(\varphi_{ts})]$ for all $s\in[0,1]$. The map $\gamma:[0,1]\to\mathbb{R}$, $s\mapsto c(\widetilde{\psi}\cdot\widetilde{\varphi}_{s})$ is continuous (by Definition 2.1) and takes value in $\mathrm{Spec}^{\alpha}(\widetilde{\psi})$ thanks to Lemma 5.6. Since $\mathrm{Spec}^{\alpha}(\widetilde{\psi})$ is nowhere dense (see [1, Lemma 2.11]) we conclude that $\gamma$ is constant and in particular $\gamma(0)=c(\widetilde{\psi})=\gamma(1)=c(\widetilde{\psi}\cdot\widetilde{\varphi})$. ∎
We are now ready to prove Theorem 5.3:
Proof of Theorem 5.3.
Consider $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})$ then
$\lceil c_{k}(\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}=\lceil c_{k}(\widetilde{\psi}^{-1}\cdot\widetilde{\psi}\cdot\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}\leq c_{k}(\widetilde{\psi}^{-1})+\lceil c_{k}(\widetilde{\psi}\cdot\widetilde{\varphi})\rceil_{\frac{2\pi}{k}}=c_{k}(\widetilde{\psi}^{-1})+\lceil c_{k}(\widetilde{\psi})\rceil_{\frac{2\pi}{k}},$ (2)
where the last equality follows from Proposition 5.5. This proves the first statement. Moreover under the assumption that $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(L_{k}^{2n+1},\xi_{k})$ one can use Proposition 3.3 to show that the relations in (2) remain true without taking the $\frac{2\pi}{k}$-ceiling function which proves the second statement. ∎
6. Estimation of the contact capacity
Let us recall our conventions to describe Hamiltonian dynamics. $(W,\Omega)$ being a symplectic manifold, to any compactly supported function $H:[0,1]\times W\to\mathbb{R}$ we associate the path of symplectic vector fields $(X_{H}^{t})$ on $W$, i.e. the flow $(\psi_{H}^{t})$ of which is an isotopy of symplectomorphisms, defined by
$\Omega(X_{t},\cdot)=-\mathrm{d}H(t,\cdot)\quad\text{ for all }t\in[0,1].$ (3)
We say that $H$ is the Hamiltonian function that generates the isotopy $(\psi_{H}^{t})$ and denote by $\mathrm{Ham}_{c}(W,\omega)$ the set of diffeomorphisms arising as time $1$-map of such isotopies.
Similarly if $(M,\xi=\ker\alpha)$ is a closed contact manifold, with a fixed contact form $\alpha$, to any function $h:[0,1]\times M\to\mathbb{R}$ one associates a path of contact vector fields $(Y_{h}^{t})$, i.e. the flow $(\phi_{h}^{t})$ of which is a contact isotopy, defined by the two equations
$\left\{\begin{array}[]{ll}\alpha(Y_{h}^{t})=h(t,\cdot)\\ \mathrm{d}\alpha(Y_{h}^{t},\cdot)=\mathrm{d}h^{t}(R_{\alpha})\alpha-\mathrm{d}h^{t}\end{array}\right.$ (4)
where $R_{\alpha}$ denotes the Reeb vector field associated to $\alpha$ (the vector field defined by the above equations (4) for the function constantly equal to $1$). We say that $h$ is the $\alpha$-Hamiltonian function that generates the contact isotopy $(\phi_{h}^{t})$.
We leave the proof of the following Lemma to the reader.
Lemma 6.1.
Let $T>0$ be any positive real number and $\pi$ be the projection of a $S^{1}:=\mathbb{R}/T$ principal bundle $M$ over a symplectic manifold $(W,\Omega)$ so that $\pi^{*}\Omega=\mathrm{d}\alpha$ for some contact form $\alpha$ on $M$, whose Reeb flow induces the $S^{1}$-action. Then for any function $H:[0,1]\times W\to\mathbb{R}$ the $\alpha$-Hamiltonian function
$h:[0,1]\times M\to\mathbb{R}\quad\quad\quad(t,x)\mapsto H(t,\pi(x))$
generates a contact isotopy $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(M,\ker\alpha)$ which is $S^{1}$-equivariant, i.e. $\phi_{h}^{t}\circ\phi_{\alpha}^{s}=\phi_{\alpha}^{s}\circ\phi_{h}^{t}$ for all $s\in\mathbb{R}$, $t\in[0,1]$. Moreover $\pi\circ\phi_{h}^{t}=\psi_{H}^{t}\circ\pi$ for all $t\in[0,1]$, where $(\psi_{H}^{t})\subset\mathrm{Ham}(W,\Omega)$ is the isotopy generated by $H$.
6.1. Estimation of the lower bound
The aim of this subsection is to prove the lower bound in Theorem 4.1 satisfied by $C_{k}$, for any integer $k\geq 2$ and $n\in\mathbb{N}_{>0}$, which we recall in the next Proposition.
Proposition 6.2.
Let $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ be a symplectic embedding where $R>0$. Then $C_{k}(B_{k}(r))\geq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
The proof of Proposition 6.2 will be an adaptation of the proof of [26, Lemma 3, Chapter 3] stating that the Hofer-Zehnder capacity of $(\mathbb{B}^{2n}(r),\omega_{0})$ is bounded from below by $\pi r^{2}$.
Proof.
Step 1 : Construction in $(\mathbb{B}^{2n}(r),\omega_{0})$
Consider $0<\varepsilon<\pi r^{2}$ and a smooth function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ supported in $[0,r^{2})$ that satisfies moreover :
  1. (1)
    $f(t)=\pi r^{2}-\varepsilon$ for $t$ in a non-empty neighborhood of $\{0\}$
  2. (2)
    $-\pi<f^{\prime}(t)\leq 0$ for all $t\in\mathbb{R}_{\geq 0}$
  3. (3)
    $f^{\prime}(t)=0$ if and only if $f(t)\in\{\pi r^{2}-\varepsilon,0\}$
Figure 1: Graphs of different functions $f$ for different values of $\varepsilon$ with $0<\varepsilon_{3}<\varepsilon_{2}<\varepsilon_{1}$
[…]
$X_{H}(x)=J\nabla H(x)=2f^{\prime}(\|x\|^{2})Jx$
where $J$ is the standard complex structure of $\mathbb{R}^{2n}$777after identifying $\mathbb{R}^{2n}$ with $\mathbb{C}^{n}$ via $\Phi$ defined in the introduction $J:=\Phi^{-1}\circ i\circ\Phi$ corresponds to the multiplication by $i$ and $\nabla H$ denotes the gradient of $H$ with respect to the standard Euclidean scalar product. A direct computation shows that for all $t\in[0,1]$ and $x\in\mathbb{B}^{2n}(r)$
$\psi_{H}^{t}(x)=e^{2f^{\prime}(\|x\|^{2})Jt}x=\cos(2f^{\prime}(\|x\|^{2})t)x+\sin(2f^{\prime}(\|x\|^{2})t)Jx,$ (5)
where $(\psi_{H}^{t})\subset\mathrm{Ham}_{c}(\mathbb{B}^{2n}(r))$ denotes the flow generated by $H$.
Note that since $2f^{\prime}([0,+\infty[)\subset(-2\pi,0]$ and $f^{\prime}(0)=0$, the orbit $t\mapsto\psi_{H}^{t}(x)$ of a point $x\in\mathbb{B}^{2n}(r)$ is $T$-periodic for some $T\in(0,1]$, i.e. $\psi_{H}^{T}(x)=x$, if and only if $f^{\prime}(\|x\|^{2})=0$. In other words the orbit of $x$ is $T$-periodic, with $T\in(0,1]$, if and only if $f(\|x\|^{2})=H(x)\in\{\pi r^{2}-\varepsilon,0\}$ and $\mathrm{d}_{x}H=0$.
Step 2 : Import the construction to $(\mathbb{C}\mathrm{P}^{n},\omega)$ Consider now on $(\mathbb{C}\mathrm{P}^{n},\omega)$ the Hamiltonian function
$G:\mathbb{C}\mathrm{P}^{n}\to\mathbb{R}\quad\quad x\mapsto\left\{\begin{array}[]{ll}H(\Psi^{-1}(x))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(r))\\ 0&\text{ otherwise}\end{array}.\right.$
A direct computation shows that the isotopy it generates $(\psi_{G}^{t})\subset\mathrm{Ham}(\mathbb{C}\mathrm{P}^{n},\omega)$ satisfies :
$\psi_{G}^{t}:x\mapsto\left\{\begin{array}[]{ll}\Psi(\psi_{H}^{t}(\Psi^{-1}(x)))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(r))\\ x&\text{ otherwise}\end{array}.\right.$
In particular, thanks to the first step we get that the orbit $t\mapsto\psi_{G}^{t}(x)$ of a point $x\in\mathbb{C}\mathrm{P}^{n}$ is $T$-periodic, with $T\in(0,1]$, if and only if $G(x)\in\{\pi r^{2}-\varepsilon,0\}$ and $\mathrm{d}_{x}G=0$.
Step 3 : Import the construction to $(L_{k}^{2n+1},\xi_{k})$ Consider on $(L_{k}^{2n+1},\ker\alpha_{k})$ the $\alpha_{k}$-Hamiltonian function
$h:L_{k}^{2n+1}\to\mathbb{R},\quad\quad x\mapsto G(\pi_{k}(x)),$
and denote by $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ the flow of contactomorphisms it generates. It is clear that its lift $(\overline{\phi_{h}^{t}})\subset\mathrm{Cont}_{0}(\mathbb{S}^{2n+1},\ker\alpha)$, starting at the identity, is generated by the $\alpha$-contact Hamiltonian $\overline{h}:\mathbb{S}^{2n+1}\to\mathbb{R}$, $x\mapsto G(\pi_{1}(x))$, where for simplicity $\alpha:=\alpha_{1}$. We claim that: if $x\in\mathbb{S}^{2n+1}$ is a $\alpha$-translated point of $\overline{\phi_{h}^{T}}$ for some $T\in(0,1]$ then
there is $\text{there is }a\in\{\pi r^{2}-\varepsilon,0\}\text{ so that }\overline{\phi_{h}^{t}}(x)=\phi_{\alpha}^{at}(x)\text{ for all }t\in[0,1].$ (6)
To prove this claim, fix $x\in\mathbb{S}^{2n+1}$ and $T\in(0,1]$ so that $x$ is a $\alpha$-translated of $\overline{\phi_{h}^{T}}$. By Lemma 6.1 $x$ is a $\alpha$-translated point of $\overline{\phi_{h}^{T}}$ implies that $\pi_{1}(x)$ is a fixed point of $\psi_{G}^{T}$. This implies, by the previous step, that $G(\pi_{1}(x))\in\{\pi r^{2}-\varepsilon,0\}$ and $\mathrm{d}_{\pi_{1}(x)}G=0$. Therefore $\mathrm{d}_{x}\overline{h}=0$ and $\overline{h}(x)\in\{\pi r^{2}-\varepsilon,0\}$. Since $\overline{h}\circ\phi_{\alpha}^{t}=\overline{h}$ for all $t\in\mathbb{R}$ we also get
$\mathrm{d}\overline{h}(R_{\alpha})=0,\quad\mathrm{d}_{\phi_{\alpha}^{t}(x)}\overline{h}=0,\quad\text{ and }\quad\overline{h}(\phi_{\alpha}^{t}(x))=\overline{h}(x).$ (7)
The last three equations (7) together with equation (4) imply that the contact vector field $Y$ associated to the $\alpha$-Hamiltonian function $\overline{h}$ satisfies for all $t\in\mathbb{R}$
$Y(\phi_{\alpha}^{t}(x))=\overline{h}(x)R_{\alpha}(\phi_{\alpha}^{t}(x)).$
Therefore the map $[0,1]\to\mathbb{S}^{2n+1}$, $t\mapsto\phi_{\alpha}^{\overline{h}(x)t}(x)$ solves the differential equation $\gamma^{\prime}(t)=Y(\gamma(t))$ with initial condition $\gamma(0)=x$. By unicity of solutions of such equations $\overline{\phi_{h}^{t}}(x)=\phi_{\alpha}^{\overline{h}(x)t}(x)$ for all $t\in[0,1]$ and the claim (6) follows. In particular we have proved that for all $T\in(0,1]$
$\mathrm{Spec}^{\alpha}(\overline{\phi_{h}^{T}})=\left\{T(\pi r^{2}-\varepsilon)+2\pi m\ |\ m\in\mathbb{Z}\right\}\bigcup\left\{2\pi m\ |\ m\in\mathbb{Z}\right\}.$ (8)
Step 4 : Computation of the selector and conclusion Let us denote by $\widetilde{\phi_{h}^{T}}:=[(\phi_{h}^{tT})]\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$ for all $T\in[0,1]$. Proposition 3.2 ensures that $c_{k}(\widetilde{\phi_{h}^{0}})=0$ and that the map $[0,1]\to\mathbb{R},\ T\mapsto c_{k}(\widetilde{\phi_{h}^{T}})\in\overline{\mathrm{Spec}^{\alpha_{k}}}(\widetilde{\phi_{h}^{T}})=\mathrm{Spec}^{\alpha}(\overline{\phi_{h}^{T}})$ is continuous. Therefore there is $a\in\{T(\pi r^{2}-\varepsilon),0\}$ so that $c_{k}(\widetilde{\phi_{h}^{T}})=Ta$ for all $T\in[0,\delta]$ where $\delta:=\min\{1,\frac{2\pi}{\pi r^{2}-\varepsilon}\}$.
Figure 2: The graphs of $f_{j}:T\mapsto T(\pi r^{2}-\varepsilon)+2\pi j$ and $g_{j}:T\mapsto 2\pi j$ for $j\in\{-1,0,1\}$
Since $\mathrm{id}\preceq\widetilde{\phi_{h}^{T}}$ and $\widetilde{\phi_{h}^{T}}\neq\mathrm{id}$, for $T>0$, we use Proposition 3.2 again to get that $c_{k}(\widetilde{\phi_{H}^{T}})>0$ and thus
$c_{k}(\widetilde{\phi_{h}^{T}})=T(\pi r^{2}-\varepsilon)\text{ for all }T\in[0,\delta].$ (9)
The relation (9) actually holds for all $T\in[0,1]$ since by [27, Thm 1] the radius $r\leq\sqrt{2}$ and so $\delta=1$ (see also Remark 1.2). The $\alpha_{k}$-Hamiltonian function $h$ being supported in $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n},(r)))$ we get by definition that
$C_{k}(B_{k}(r))\geq c_{k}(\widetilde{\phi_{h}^{1}})=\pi r^{2}-\varepsilon.$
Since $\varepsilon>0$ can be choosen as small as we want this finishes the proof. ∎
6.2. Estimation of the upper bound
The aim of this subsection is to prove the upper bound in Theorem 4.1, for all $n\in\mathbb{N}_{>0}$ and integer $k\geq 2$, which we recall in the following Proposition.
Proposition 6.3.
Consider $\Psi:(\mathbb{B}^{2n}(\lambda R),\omega_{0})\hookrightarrow(\mathbb{C}\mathrm{P}^{n},\omega)$ a symplectic embedding where $R>0$ and $\lambda>\sqrt{3}$. Then for all $r\in(0,R]$
$C_{k}(B_{k}(r))\leq\pi r^{2}$
where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
Let us first consider $(\mathbb{R}^{2}=\mathbb{R}^{2}\times\{0\},dx_{1}\wedge dy_{1})$ the symplectic submanifold of $(\mathbb{R}^{2n},\sum\limits_{i=1}^{n}dx_{i}\wedge dy_{i})$. We recall a Lemma from [25, Lemma 4.2].
Lemma 6.4 ([25]).
For every $\delta>0$ there exists an Hamiltonian function $H:\mathbb{R}^{2}\to\mathbb{R}_{\geq 0}$ supported in $\mathbb{B}^{2}(\sqrt{2+2\delta})$ so that $\max H\leq(\pi+\delta)$ and $\psi_{H}^{1}(\mathbb{B}^{2}(1))\bigcap\mathbb{B}^{2}(1)=\emptyset$.
As a direct corollary we have.
Corollary 6.5.
For every $r,\delta>0$ there exists an Hamitlonian function $H:\mathbb{R}^{2}\to\mathbb{R}_{\geq 0}$ supported in $\mathbb{B}^{2}(r\sqrt{2+2\delta})$ so that $\max H\leq r^{2}(\pi+\delta)$ and $\psi_{H}^{1}(\mathbb{B}^{2}(r))\bigcap\mathbb{B}^{2}(r)=\emptyset$.
Proof.
Let $H$ be a Hamiltonian function that satisfies the requirements of Lemma 6.4 with respect to $\delta>0$ and consider the diffeomorphism $\psi_{r}:\mathbb{R}^{2}\to\mathbb{R}^{2},\ x\mapsto rx$. A direct computation shows that the isotopy $(\psi_{r}\circ\psi_{t}\circ\psi_{r}^{-1})$ is contained in $\mathrm{Ham}_{c}(\mathbb{R}^{2},\omega_{0})$ and is generated by the Hamiltonian function $H_{r}:=r^{2}H\circ(\psi_{r})^{-1}$ whose support is equal to $(\psi_{r})(\operatorname{supp}(H))$. Therefore $H_{r}$ satisfies the requirements of Corollary 6.5 with respect to $\delta>0$ and $r>0$. ∎
We are now ready to prove Proposition 6.3.
Proof of Proposition 6.3.
Step 1 : Construction in $(\mathbb{B}^{2n}(\lambda R),\omega_{0})$
Let $r\in(0,R]$, $\delta>0$ be small enough so that $\sqrt{3+2\delta}<\lambda$ and $H:\mathbb{R}^{2}\to\mathbb{R}_{\geq 0}$ be an Hamiltonian function supported in $\mathbb{B}^{2}(r\sqrt{2+2\delta})$ so that $\max H\leq r^{2}(\pi+\delta)$ and $\psi_{H}^{1}(\mathbb{B}^{2}(r))\bigcap\mathbb{B}^{2}(r)=\emptyset$ whose existence is guaranteed by Corollary 6.5. The function $K:\mathbb{R}^{2n}\to\mathbb{R}_{\geq 0},((x_{1},y_{1}),\cdots,(x_{n},y_{n}))\mapsto H(x_{1},y_{1})$, even if non compactly supported, generates a vector field $X_{k}$, i.e. $\omega_{0}(X_{k},\cdot)=-dK$, which is complete, and whose flow $(\psi_{K}^{t})\subset\mathrm{Ham}(\mathbb{R}^{2n},\omega_{0})$ satisfies $\psi_{K}^{t}=\psi_{H}^{t}\times\mathrm{id}_{\mathbb{R}^{2(n-1)}}$ for all $t\in[0,1]$. Therefore
$\psi_{K}^{1}\left(\mathbb{B}^{2n}(r)\right)\bigcap\mathbb{B}^{2n}(r)\subset\left(\psi_{H}^{1}\left(\mathbb{B}^{2}(r)\right)\times\mathbb{B}^{2(n-1)}(r)\right)\bigcap\mathbb{B}^{2}(r)\times\mathbb{B}^{2(n-1)}(r)=\emptyset.$
Moreover $\underset{t\in[0,1]}{\bigcup}\psi_{K}^{t}(\mathbb{B}^{2n}(r))$ is contained in $\mathbb{B}^{2n}(r\sqrt{3+2\delta})$ since
$\psi_{K}^{t}(\mathbb{B}^{2n}(r))\subset\psi_{H}^{t}(\mathbb{B}^{2}(r))\times\mathbb{B}^{2(n-1)}(r)\subset\mathbb{B}^{2}(r\sqrt{2+2\delta})\times\mathbb{B}^{2(n-1)}(r)$
and the latter set is clearly contained inside $\mathbb{B}^{2n}(r\sqrt{3+2\delta})$. Given that $\sqrt{3+2\delta}<\lambda$ one can construct $\rho:\mathbb{R}^{2n}\to[0,1]$ a smooth cutoff function supported in $\mathbb{B}^{2n}(\lambda R)$ which is constantly equal to $1$ on $\mathbb{B}^{2n}(r\sqrt{3+2\delta})$ so that the Hamiltonian function $\widetilde{K}:=\rho K$ generates a path of Hamiltonian symplectomorphisms $(\psi_{t})$ satisfying
  1. (1)
    $\widetilde{K}$ and $(\psi_{t})$ are both supported in $\mathbb{B}^{2n}(\lambda R)$
  2. (2)
    $\widetilde{K}\geq 0$ and $\max\widetilde{K}\leq r^{2}(\pi+\delta)$
  3. (3)
    $\psi_{1}\left(\mathbb{B}^{2n}(r)\right)\cap\mathbb{B}^{2n}(r)=\emptyset$.
Step 2 : Import the construction to $(\mathbb{C}\mathrm{P}^{n},\omega)$ A direct computation shows that the Hamiltonian function
$G:\mathbb{C}\mathrm{P}^{n}\to\mathbb{R}_{\geq 0},\quad\quad x\mapsto\left\{\begin{array}[]{ll}\widetilde{K}(\Psi^{-1}(x))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(\lambda R))\\ 0&\text{ otherwise}\end{array}\right.$
on $(\mathbb{C}\mathrm{P}^{n},\omega)$, generates an isotopy $(\psi_{G}^{t})\subset\mathrm{Ham}(\mathbb{C}\mathrm{P}^{n},\omega)$ that satisfies
$[0,1]\times\mathbb{C}\mathrm{P}^{n}\to\mathbb{C}\mathrm{P}^{n},\quad\quad(t,x)\mapsto\left\{\begin{array}[]{ll}\Psi(\psi_{t}(\Psi^{-1}(x)))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(\lambda R))\\ x&\text{ otherwise.}\end{array}\right.$
We therefore clearly have:
  1. (1)
    $G\geq 0$ and $\max G\leq r^{2}(\pi+\delta)$
  2. (2)
    $\psi_{G}^{1}(B(r))\cap B(r)=\emptyset$, where $B(r):=\Psi(\mathbb{B}^{2n}(r))$.
Step 3 : Lift the construction to $(L_{k}^{2n+1},\xi_{k})$ We lift this construction to $(L_{k}^{2n+1},\ker\alpha_{k})$, thanks to Lemma 6.1. More precisely the $\alpha_{k}$-Hamiltonian function $h:L_{k}^{2n+1}\to\mathbb{R}_{\geq 0}$, $x\mapsto G(\pi_{k}(x))$ generates a flow of $\mathbb{S}^{1}$-equivariant contactomorphisms $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(L_{k}^{2n+1},\xi_{k})$ satisfying :
  1. (1)
    $h\geq 0$ and $\max h\leq r^{2}(\pi+\delta)$
  2. (2)
    $\phi_{h}^{1}(B_{k}(r))\cap B_{k}(r)=\emptyset$, where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
Indeed, if we suppose by contradiction that $\phi_{h}^{1}(B_{k}(r))\cap B_{k}(r)\neq\emptyset$ then we get the following absurdity
$\emptyset\neq\pi_{k}\left(\phi_{h}^{1}(B_{k}(r))\cap B_{k}(r)\right)\subset\pi_{k}\left(\phi_{h}^{1}(B_{k}(r))\right)\cap\pi_{k}\left(B_{k}(r)\right)=\psi_{G}^{1}(B(r))\cap B(r)=\emptyset.$
Step 4 : Conclusion Therefore thanks to Theorem 5.3
$c_{k}(B_{k}(r))\leq c_{k}(\widetilde{\phi}^{-1})+c_{k}(\widetilde{\phi})$
where $\widetilde{\phi}:=[(\phi_{h}^{t})]\in\widetilde{\mathrm{Cont}_{0}}(L_{k}^{2n+1},\xi_{k})$. Finally, since the $\alpha_{k}$-Hamiltonian function $L_{k}^{2n+1}\to\mathbb{R}_{\leq 0},\ x\mapsto-h(x)$, generates the contact isotopy $(\phi_{h}^{t})^{-1}$ which represents $\widetilde{\phi}^{-1}$, the last point of Proposition 3.2 implies that
$c_{k}(\widetilde{\phi})+c_{k}(\widetilde{\phi}^{-1})\leq\int_{0}^{1}\max h\mathrm{d}t-\int_{0}^{1}\min h\mathrm{d}t\leq r^{2}(\pi+\delta).$
Letting $\delta$ goes to 00 finishes the proof. ∎
7. Order spectral selectors for strongly orderable prequantizations
7.1. Order spectral selectors on Legendrian isotopy class
Let $(M,\xi)$ be a cooriented contact manifold, not necessarily closed. As in the closed case in Section 2, a contact isotopy, that we denote by $(\phi_{t})$, is a $[0,1]$-family of contactomorphisms $\{\phi_{t}\}\subset\mathrm{Cont}(M,\xi)$ so that $[0,1]\times M\to M$ is smooth, and we denote by $\mathrm{Cont}_{0}(M,\xi)$ the set of contactomorphisms isotopic to the identity, and by $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ the ones that are isotopic to the identity through a compactly supported isotopy. Similarly, a $[0,1]$-family of closed Legendrians $(\Lambda_{t})\subset(M,\xi)$ is a Legendrian isotopy if and only if there exists a compactly supported contact isotopy $(\phi_{t})\subset\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$ parametrizing it, i.e. $\phi_{t}(\Lambda_{0})=\Lambda_{t}$ and $\phi_{t}$ is compactly supported. Note that all the Legendrians we deal with in this paper are closed (even if the contact manifold itself may not). Therefore the isotopy class of a closed Legendrian $\Lambda\subset M$, that we denote by $\mathcal{L}(\Lambda)$, is an homogeneous space $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)/\mathrm{Stab}(\Lambda)$ and inherits the $C^{1}$-topology of $\mathrm{Cont}_{0}^{\mathrm{c}}(M,\xi)$. We denote by ~$\widetilde{\mathcal{L}}(\Lambda)$ its universal cover and $\Pi:\widetilde{\mathcal{L}}(\Lambda)\to\mathcal{L}(\Lambda)$ the natural projection. Note that we have a natural left action of $\mathrm{Cont}_{0}(M,\xi)$ (resp. $\widetilde{\mathrm{Cont}_{0}}(M,\xi)$888the set of equivalence classes of contact isotopy starting at the identity, where two such contact isotopies are identified if they are homotopic relatively to endpoints) on $\mathcal{L}(\Lambda)$ (resp. ~$\widetilde{\mathcal{L}}(\Lambda)$).
If $\alpha$ is a contact form whose Reeb vector field is complete, then as for contactomorphisms we have a notion of spectrum. More precisely let $\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0}\in\widetilde{\mathcal{L}}(\Lambda)$ and denote by $\Lambda_{1}$ and $\Lambda_{0}$ their respective projections to $\mathcal{L}(\Lambda)$ then
$\mathrm{Spec}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})=\{t\in\mathbb{R}\ |\ \phi_{\alpha}^{t}(\Lambda_{0})\cap\Lambda_{1}\neq\emptyset\}.$
A Legendrian isotopy $(\Lambda_{t})$ is said to be non-negative (resp. positive) if for some, and thus all, contact form $\alpha$ supporting $\xi$ we have $\alpha(\frac{d}{dt}\phi_{t}(x))\geq 0$ (resp. $>0$) for all $x\in\Lambda_{0}$ and any compactly supported contact isotopy $(\phi_{t})$ parametrizing it. We say, following [16], that ~$\widetilde{\mathcal{L}}(\Lambda)$ is orderable if there does not exist any positive contractible loop of Legendrians in $\mathcal{L}(\Lambda)$. Assuming that ~$\widetilde{\mathcal{L}}(\Lambda)$ is orderable implies that the binary relation $\preceq$, defined by $\widetilde{\Lambda}_{0}\preceq\widetilde{\Lambda}_{1}$ if and only if there exists a path $(\widetilde{\Lambda}_{t})$ joining $\widetilde{\Lambda}_{0}$ to $\widetilde{\Lambda}_{1}$ such that $(\Pi(\widetilde{\Lambda}_{t}))$ is a non-negative Legendrian isotopy, is a partial order.
For any contact form $\alpha$ supporting $\xi$ whose Reeb vector field is complete in [3] the authors defined a map $\ell_{+}^{\alpha}:\widetilde{\mathcal{L}}(\Lambda)\times\widetilde{\mathcal{L}}(\Lambda)\to\mathbb{R}\cup\{-\infty\}$
$(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})\mapsto\inf\left\{t\in\mathbb{R}\ |\widetilde{\Lambda}_{1}\preceq\widetilde{\phi_{\alpha}^{t}}\cdot\widetilde{\Lambda}_{0}\right\}=\inf\left\{\underset{[0,1]\times\Lambda_{0}}{\max}\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\ |\ [(\phi_{t})]\cdot\widetilde{\Lambda}_{0}=\widetilde{\Lambda}_{1}\right\},$
where $\Lambda_{0}:=\Pi(\widetilde{\Lambda}_{0})$, and showed the following:
Proposition 7.1 ([3]).
The map $\ell_{+}^{\alpha}$ takes value in $\mathbb{R}$ if and only if ~$\widetilde{\mathcal{L}}(\Lambda)$ is orderable. Moreover if ~$\widetilde{\mathcal{L}}(\Lambda)$ is orderable then for any $\widetilde{\Lambda}_{0},\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{2}\in\widetilde{\mathcal{L}}(\Lambda)$ and any $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$
  1. (1)
    $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{0},\widetilde{\Lambda}_{0})=0$
  2. (2)
    $\widetilde{\Lambda}_{0}\preceq\widetilde{\Lambda}_{1}$ implies that $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})\geq 0$ and $\ell_{+}^{\alpha}(\Lambda_{0},\Lambda_{1})\leq 0$
  3. (3)
    $\widetilde{\Lambda}_{0}\preceq\widetilde{\Lambda}_{1}$ and $\widetilde{\Lambda}_{0}\neq\widetilde{\Lambda}_{1}$ implies that $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})>0$
  4. (4)
    $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{2},\widetilde{\Lambda}_{0})\leq\ell_{+}^{\alpha}(\widetilde{\Lambda}_{2},\widetilde{\Lambda}_{1})+\ell_{+}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})$
  5. (5)
    $\ell_{+}^{\alpha}(\widetilde{\varphi}\cdot\widetilde{\Lambda}_{1},\widetilde{\varphi}\cdot\widetilde{\Lambda}_{0})=\ell_{+}^{\varphi^{*}\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})$ where $\varphi=\Pi(\widetilde{\varphi})$
  6. (6)
    For any contact isotopy $(\phi_{t})$ representing $\widetilde{\varphi}$
    $\int_{0}^{1}\underset{x\in\Pi(\widetilde{\Lambda}_{0})}{\min}\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\mathrm{d}t\leq\ell_{+}^{\alpha}(\widetilde{\varphi}\widetilde{\Lambda}_{0},\widetilde{\Lambda}_{0})\leq\int_{0}^{1}\underset{x\in\Pi(\widetilde{\Lambda}_{0})}{\max}\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\mathrm{d}t$
  7. (7)
    $\ell_{+}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})\in\mathrm{Spec}^{\alpha}(\widetilde{\Lambda}_{1},\widetilde{\Lambda}_{0})$.
7.2. Spectral selectors for strongly orderable closed contact manifold
Let $(M,\xi=\ker\alpha)$ be a closed cooriented contact manifold and consider its contact product $(M\times M\times\mathbb{R},\Xi=\ker(\beta:=\mathrm{pr}_{2}^{*}\alpha-e^{\theta}\mathrm{pr}_{1}^{*}\alpha))$ as in Section 1.2. To any $[(\varphi_{t})]=\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ one can associate $[(\Phi_{t})]=\widetilde{\Phi}\in\widetilde{\mathrm{Cont}_{0}}(M\times M\times\mathbb{R},\Xi)$ defined by
$\Phi_{t}:(x_{1},x_{2},\theta)\mapsto(x_{1},\varphi_{t}(x_{2}),\theta+g_{t}(x_{2})),$ (10)
where $g_{t}$ is the conformal factor of $\varphi_{t}$ with respect to $\alpha$, i.e. $\phi_{t}^{*}\alpha=e^{g_{t}}\alpha$. We say that $\widetilde{\Phi}$ is the graph of $\widetilde{\varphi}$. Direct computations show that
$\mathrm{Spec}^{\alpha}(\widetilde{\varphi})=\mathrm{Spec}^{\beta}(\widetilde{\Phi}\cdot\Delta,\Delta),$ (11)
where we have identified the Legendrian $\Delta=\{(x,x,0)\ |\ x\in M\}\subset(M\times M\times\mathbb{R},\ker\beta)$ with the equivalence class of the constant isotopy at $\Delta$ in ~$\widetilde{\mathcal{L}}(\Delta)$, and that for all $t\in[0,1]$ and $(x_{1},x_{2},\theta)\in M\times M\times\mathbb{R}$
$\beta\left(\frac{d}{dt}\Phi_{t}(x_{1},x_{2},\theta)\right)=\alpha\left(\frac{d}{dt}\phi_{t}(x_{2})\right).$ (12)
Suppose that in addition to be closed $(M,\xi=\ker\alpha)$ is also strongly orderable, i.e. ~$\widetilde{\mathcal{L}}(\Delta)$ is orderable, then using Proposition 7.1 we can define the map
$C_{+}^{\alpha}:\widetilde{\mathrm{Cont}_{0}}(M,\xi)\to\mathbb{R}\quad\quad\widetilde{\varphi}\mapsto\ell_{+}^{\beta}(\widetilde{\Phi}\cdot\Delta,\Delta),\text{ where }\widetilde{\Phi}\text{ is the graph of }\widetilde{\varphi}$
which satisfies moreover the following:
Proposition 7.2.
Suppose that $(M,\xi=\ker\alpha)$ is a closed strongly orderable contact manifold. Then for any $\widetilde{\varphi},\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\xi)$ denoting by $\widetilde{\Phi},\widetilde{\Psi}\in\widetilde{\mathrm{Cont}_{0}}(M\times M\times\mathbb{R},\ker\beta)$ their corresponding graphs
  1. (1)
    $C_{+}^{\alpha}(\mathrm{id})=0$
  2. (2)
    $\widetilde{\psi}\preceq\widetilde{\varphi}$ implies that $C_{+}^{\alpha}(\widetilde{\psi})\leq C_{+}^{\alpha}(\widetilde{\varphi})$
  3. (3)
    $\mathrm{id}\preceq\widetilde{\varphi}$ and $\widetilde{\varphi}\neq\mathrm{id}$ implies that $C_{+}^{\alpha}(\widetilde{\varphi})>0$
  4. (4)
    $C_{+}^{\alpha}(\widetilde{\psi}\widetilde{\varphi})\leq\ell_{+}^{\Psi^{*}\beta}(\widetilde{\Phi}\Delta,\Delta)+C_{+}^{\alpha}(\widetilde{\psi})$ where $\Psi:=\Pi(\widetilde{\Psi})\in\mathrm{Cont}_{0}(M\times M\times\mathbb{R},\Xi)$
  5. (5)
    For any contact isotopy $(\phi_{t})$ representing $\widetilde{\varphi}$
    $\int_{0}^{1}\min\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\mathrm{d}t\leq C_{+}^{\alpha}(\widetilde{\varphi})\leq\int_{0}^{1}\max\alpha\left(\frac{d}{dt}\phi_{t}(x)\right)\mathrm{d}t$
  6. (6)
    $C_{+}^{\alpha}(\widetilde{\varphi})\in\mathrm{Spec}^{\alpha}(\widetilde{\varphi})$.
Proof.
  1. (1)
    $C_{+}^{\alpha}(\mathrm{id})=\ell_{+}^{\beta}(\Delta,\Delta)=0$
  2. (2)
    By the triangle inequality of Proposition 7.1 we have
    $\ell_{+}^{\beta}(\widetilde{\Phi}\cdot\Delta,\Delta)-\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\Delta)$ $\geq\ell_{+}^{\beta}(\widetilde{\Phi}(\Delta),\Delta)-\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\widetilde{\Phi}\cdot\Delta)-\ell_{+}^{\beta}(\widetilde{\Phi}\cdot\Delta,\Delta)$
    $=-\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\widetilde{\Phi}\cdot\Delta).$
    If we suppose that $\widetilde{\psi}\preceq\widetilde{\varphi}$ this implies, using (12), that $\widetilde{\Psi}\cdot\Lambda\preceq\widetilde{\Phi}\cdot\Lambda$ and thus using the monotonicity property of $\ell_{+}^{\beta}$ we get the desired result
    $C_{+}^{\alpha}(\widetilde{\varphi})-C_{+}^{\alpha}(\widetilde{\psi})\geq-\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\widetilde{\Phi}\cdot\Delta)\geq 0.$
  3. (3)
    Let $\widetilde{\varphi}$ be such that $\mathrm{id}\preceq\widetilde{\varphi}$ and $\widetilde{\varphi}\neq\mathrm{id}$. Let $(\phi_{t})$ be a non-negative contact isotopy so that $[(\phi_{t})]=\widetilde{\varphi}$ and consider its graph $(\Phi_{t})$ which is also non-negative contact isotopy by (12). Since $\widetilde{\varphi}\neq\mathrm{id}$ there exists $T\in[0,1]$ and $x\in M$ so that $\phi_{T}(x)\neq x$ which implies that $\Phi_{T}(\Delta)\neq\Delta$. Therefore using the non-degeneracy and monotonicity of $\ell_{+}^{\beta}$ we get that
    $C_{+}^{\alpha}(\widetilde{\varphi})=\ell_{+}^{\beta}(\widetilde{\Phi}_{1}\cdot\Delta,\Delta)\geq\ell_{+}^{\beta}(\widetilde{\Phi}_{T}\cdot\Delta,\Delta)>0$
  4. (4)
    By the triangle inequality property and the naturality of $\ell_{+}^{\beta}$ we get that
    $\ell_{+}^{\beta}(\widetilde{\Psi}\widetilde{\Phi}\cdot\Delta,\Delta)\leq\ell_{+}^{\beta}(\widetilde{\Psi}\widetilde{\Phi}\cdot\Delta,\widetilde{\Psi}\cdot\Delta)+\ell_{+}^{\beta}(\widetilde{\Psi}\cdot\Delta,\Delta)=\ell_{+}^{\Psi^{*}\beta}(\widetilde{\Phi}\cdot\Delta,\Delta)+C_{+}^{\alpha}(\widetilde{\psi}).$
  5. (5)
    The estimates come from the one of Proposition 7.1 combined with (12).
  6. (6)
    The spectrality of $C_{+}^{\alpha}$ follows from (11) and the spectrality of $\ell_{+}^{\beta}$.
As a corollary we get the following.
Corollary 7.3.
Suppose that $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha)$ can be represented by a path $(\psi_{t})$ so that $\psi_{t}^{*}\alpha=\alpha$ for all $t\in[0,1]$ then
$C_{+}^{\alpha}(\widetilde{\psi}\widetilde{\varphi})\leq C_{+}^{\alpha}(\widetilde{\psi})+C_{+}^{\alpha}(\widetilde{\varphi}).$
Proof.
Fist note that for any $\psi\in\mathrm{Cont}_{0}(M,\ker\alpha)$ if $g:M\to\mathbb{R}$ is the conformal factor of $\psi$ then the conformal factor of its graph $\Psi\in\mathrm{Cont}_{0}(M\times M\times\mathbb{R},\ker\beta)$ is given by $G:M\times M\times\mathbb{R}\to\mathbb{R},\ (x_{1},x_{2},\theta)\mapsto g(x_{2})$. Therefore if $\psi^{*}\alpha=\alpha$ it implies that $\Psi^{*}\beta=\beta$. The proof of the corollary follows now directly from the triangle inequality of Proposition 7.2. ∎
Remark 7.4.
If $\psi\in\mathrm{Cont}_{0}(M,\xi)$ is such that $\psi^{*}\alpha=\alpha$ then $\psi\circ\phi_{\alpha}^{t}=\phi_{\alpha}^{t}\circ\psi$ for all $t\in\mathbb{R}$. Indeed, it implies that $\psi^{*}R_{\alpha}=R_{\alpha}$
7.3. Order spectral selectors for strongly orderable prequantizations
Let $(M,\ker\alpha)$ be a cooriented contact manifold and $\Gamma$ a finite group acting freely on $M$ by $\alpha$-strict contactomorphism, i.e. $\gamma^{*}\alpha=\alpha$ for any $\gamma\in G$. Then $\alpha$ descends to a contact form $\alpha^{\prime}$ on $M^{\prime}:=M/\Gamma$ and the subgroup $\mathrm{Cont}_{0}^{\Gamma}(M,\ker\alpha)\subset\mathrm{Cont}_{0}(M,\ker\alpha)$ of contactomorphisms isotopic to the identity that are equivarient with respect to the action of $\Gamma$ is a covering space over $\mathrm{Cont}_{0}(M^{\prime},\ker\alpha^{\prime})$. In particular any contact isotopy starting at the identity in $\mathrm{Cont}_{0}(M^{\prime},\ker\alpha^{\prime})$ admits a unique lift to $\mathrm{Cont}_{0}^{\Gamma}(M,\ker\alpha)$ starting at the identity and so we deduce a lifting map
$\widetilde{\mathrm{Cont}_{0}}(M^{\prime},\ker\alpha^{\prime})\to\widetilde{\mathrm{Cont}_{0}}(M,\ker\alpha).$ (13)
Recall that a contact manifold $(M_{1},\xi_{1}^{\prime})$ is a prequantization over a symplectic manifold $(W,\omega^{\prime})$ if there exists a $\mathbb{T}^{1}$-principal bundle $\pi_{1}^{\prime}:M_{1}\to W$, where $\mathbb{T}^{1}=\mathbb{R}/2\pi\mathbb{Z}$, and a contact form $\alpha_{1}^{\prime}$ supporting $\xi_{1}^{\prime}$ so that the Reeb flow of $\alpha_{1}^{\prime}$ is Zoll of minimal period $2\pi$, induces the $\mathbb{T}^{1}$-action and satisfies $\pi_{1}^{\prime*}\omega^{\prime}=\mathrm{d}\alpha^{\prime}_{1}$.
From now on until the end of the paper, we drop all the primes from our notations, and $(M_{1},\xi_{1}=\ker\alpha_{1})$ is any closed prequantization over some symplectic manifold $(W,\omega)$ and $\alpha_{1}$ is the contact form whose Reeb flow is Zoll of minimal period $2\pi$, induces the $\mathbb{T}^{1}=\mathbb{R}/2\pi\mathbb{Z}$ action on $M_{1}$ and satisfies $\pi_{1}^{*}\omega=\alpha_{1}$, where $\pi_{1}:M_{1}\to W=M_{1}/\mathbb{T}^{1}$ is the natural projection. For any $k\in\mathbb{N}_{>0}$, $\mathbb{Z}_{k}$ is a finite group acting on $(M_{1},\xi_{1})$, by $[j]\cdot x=\phi_{\alpha_{1}}^{\frac{2\pi j}{k}}(x)$ for all $j\in\mathbb{Z}_{k},\ x\in M_{1}$, and we denote by $M_{k}:=M_{1}/\mathbb{Z}_{k}$ the quotient manifold, by $\xi_{k}$ the contact distribution given by the kernel of the induced contact $\alpha_{k}$ and by $\pi_{k}:M_{k}\to W$ the natural projection (see 1.2). We have the following Lemma.
Lemma 7.5.
If $(M_{1},\xi_{1})$ is a strongly orderable prequantization then for any $k\in\mathbb{N}_{>0}$
  1. (1)
    $(M_{k},\xi_{k})$ is strongly orderable
  2. (2)
    for any $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}}(M_{k},\xi_{k})$ denoting by $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}(M_{1},\xi_{1})$ its lift we have
    $C_{+}^{\alpha_{k}}(\widetilde{\varphi})\geq C_{+}^{\alpha_{1}}(\widetilde{\psi}).$
  3. (3)
    $\left\lceil C_{+}^{\alpha_{k}}(\widetilde{\psi}\widetilde{\varphi}\widetilde{\psi}^{-1})\right\rceil_{\frac{2\pi}{k}}=\left\lceil C_{+}^{\alpha_{k}}(\widetilde{\varphi})\right\rceil_{\frac{2\pi}{k}}$.
Proof.
In the following $(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k}=\ker\beta_{k})$ denote the contact product of $(M_{k},\xi_{k}=\ker\alpha_{k})$ with itself and $\beta_{k}$ the product contact form associated to $\alpha_{k}$ for any $k\in\mathbb{N}_{>0}$ as in Section 1.2.
  1. (1)
    Suppose that $(M_{k},\xi_{k})$ is not strongly orderable. It means that there exists a positive contractible loop of Legendrians $(\Lambda_{t})\subset(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$ based at $\Delta_{k}:=\{(x,x,0)\ |\ x\in M_{k}\}$. Consider $(\Lambda_{t}^{s})$ a homotopy of this loop with fixed endpoint to the constant loop based at $\Delta_{k}$, i.e. $\Lambda_{t}^{0}=\Lambda_{t}$ and $\Lambda_{t}^{1}=\Lambda_{1}^{s}=\Lambda_{0}^{s}=\Delta_{k}$ for all $t,\ s\in[0,1]$, and consider $(\Phi_{t}^{s})\subset\mathrm{Cont}_{0}(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$ so that $\Phi_{t}^{s}(\Delta_{k})=\Lambda_{t}^{s}$ and $\Phi_{0}^{s}=\mathrm{id}$ for all $t,s\in[0,1]$ (for instance see [3, Lemma 2.1] for the construction of $(\Phi_{t}^{s})$). Seeing $M_{k}\times M_{k}\times\mathbb{R}$ as the quotient of $M_{1}\times M_{1}\times\mathbb{R}$ by the action by strict contactomorphisms of $\mathbb{Z}_{k}\times\mathbb{Z}_{k}$, i.e. $([j_{1}],[j_{2}])\cdot(x_{1},x_{2},\theta)=([j_{1}]\cdot x_{1},[j_{2}]\cdot x_{2},\theta)$, the group $\mathrm{Cont}_{0}^{\mathbb{Z}_{k}\times\mathbb{Z}_{k}}(M_{1}\times M_{1}\times\mathbb{R},\Xi_{1})$ is a covering space over $\mathrm{Cont}_{0}(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$ and therefore there exists a unique $(\Psi_{t}^{s})\subset\mathrm{Cont}_{0}^{\mathbb{Z}_{k}\times\mathbb{Z}_{k}}(M_{1}\times M_{1}\times\Xi_{1})$ that lifts $(\Phi_{t}^{s})$, i.e. $\Pi_{k}\circ\Psi_{t}^{s}=\Phi_{t}^{s}\circ\Pi_{k}$ where $\Pi_{k}:M_{1}\times M_{1}\times\mathbb{R}\to M_{k}\times M_{k}\times\mathbb{R}$ denotes the natural projection, and so that $\Psi_{0}^{s}=\mathrm{id}$ for all $s,t\in[0,1]$. It is clear that $(\Psi_{t}^{1}(\Delta_{1}))$ is a contractible loop in $\mathcal{L}(\Delta_{1})$, and moreover since $\Pi_{k}^{*}\beta_{k}=\beta_{1}$ a direct computation shows that it is a positive loop which implies that $(M_{1},\xi_{1})$ is not strongly orderable.
  2. (2)
    Let $\widetilde{\varphi}\in\mathrm{Cont}_{0}(M_{k},\xi_{k})$ and its graph $\widetilde{\Phi}\in\mathrm{Cont}_{0}(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$. Denoting by $\widetilde{\psi}$ the lift $\widetilde{\varphi}$ to $\mathrm{Cont}_{0}(M_{1},\xi_{1})$, it is easy to see that its graph $\widetilde{\Psi}\in\mathrm{Cont}_{0}(M_{1}\times M_{1},\times\mathbb{R},\Xi_{1})$ corresponds to the $\mathbb{Z}_{k}\times\mathbb{Z}_{k}$-equivariant lift of $\widetilde{\Phi}$. Consider any contact isotopy $(\eta_{t})\subset\mathrm{Cont}_{0}(M_{k}\times M_{k}\times\mathbb{R},\Xi_{k})$ starting at the identity such that $[(\eta_{t})]\cdot\Delta_{k}=\widetilde{\Phi}\cdot\Delta_{k}$, and $(\overline{\eta}_{t})\subset\mathrm{Cont}_{0}^{\mathbb{Z}_{k}\times\mathbb{Z}_{k}}(M_{1}\times M_{1}\times\mathbb{R},\Xi_{1})$ its corresponding $\mathbb{Z}_{k}\times\mathbb{Z}_{k}$-equivariant lift starting at the identity. We claim that $[(\overline{\eta_{t}})]\cdot\Delta_{1}=\widetilde{\Psi}\cdot\Delta_{1}$. Admitting this claim for the moment finishes the proof since $\underset{p\in\Delta_{1}}{\max}\beta_{1}(\frac{d}{dt}\overline{\eta}_{t}(p))=\underset{p\in\Delta_{k}}{\max}\beta_{k}(\frac{d}{dt}\eta_{t}(p))$. To prove the claim take any two contact isotopies $(\Phi_{t})$ and $(\Psi_{t})$ representing $\widetilde{\Phi}$ and $\widetilde{\Psi}$ respectively, i.e. $[(\Phi_{t})]=\widetilde{\Phi}$ and $[(\Psi_{t})]=\widetilde{\Psi}$. Then the Legendrian isotopy $(\Pi_{k}(\Psi_{t}^{-1}\circ\overline{\eta}_{t})\Lambda_{1})=(\Phi_{t}^{-1}\circ\eta_{t}(\Lambda_{k}))$ is a contractible loop based at $\Delta_{k}$ in $\mathcal{L}(\Delta_{k})$ since $[(\eta_{t})]\cdot\Delta_{k}=\widetilde{\Phi}\cdot\Delta_{k}$. Therefore the Legendrian isotopy $((\Psi_{t}^{-1}\circ\overline{\eta}_{t})\Lambda_{1})$ has to be itself a contractible loop in $\mathcal{L}(\Delta_{1})$ which concludes the proof of our claim.
  3. (3)
    It is a direct consequence of 2.2.
8. Computation of the contact capacity in strongly orderable prequantizations
Taking the same notations as in Section 7.3 above, suppose from now on that $(M_{1},\xi_{1}=\ker\alpha_{1})$ is a closed strongly orderable prequantization over $(W,\omega)$. Thanks to Lemma 7.5 the associated contact manifold $(M_{k},\xi_{k})$ is also strongly orderable, for all $k\in\mathbb{N}_{>0}$, and therefore one can define the spectral selectors $C_{+}^{\alpha_{k}}:\widetilde{\mathrm{Cont}_{0}}(M_{k},\xi_{k})\to\mathbb{R}$ of Section 7.2). Then, as in Section 4, for any $k\in\mathbb{N}_{>0}$999here $k$ is not necessarily greater than $1$ and any non-empty open set $U\subset M_{k}$ we denote by $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})$ the subgroup of $\widetilde{\mathrm{Cont}_{0}}(M_{k},\xi_{k})$ whose elements can be represented by contact isotopies $(\phi_{t})$ supported in $U$. We define similarly $C_{\alpha_{k}}:\{\text{non-empty open sets of }M_{k}\}\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ to be the map:
$U\mapsto C_{\alpha_{k}}(U):=\sup\{C^{\alpha_{k}}_{+}(\widetilde{\varphi})\ |\ \widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})\}.$
Recall that $\pi_{k}:M_{k}\to W$ denotes the natural projection for all $k\in\mathbb{N}_{>0}$. As for lens spaces we have the following:
Theorem 8.1.
For all non-empty open sets $U,V\subset M_{k}$ and $\phi\in\mathrm{Cont}_{0}(M_{k},\xi_{k})$
  1. (1)
    (monotonicity) $C_{\alpha_{k}}(U)\leq C_{\alpha_{k}}(V)$ if $U\subset V$
  2. (2)
    (invariance by conjugation) $\lceil C_{\alpha_{k}}(U)\rceil_{\frac{2\pi}{k}}=\lceil C_{\alpha_{k}}(\phi(U))\rceil_{\frac{2\pi}{k}}$
  1. (3)
    (lower bound) if $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(W,\omega)$ is a symplectic embedding for some $R\in(0,\sqrt{2}]$ then $C_{\alpha_{k}}(B_{k}(r))\geq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$,
  2. (4)
    (upper bound) if $\Psi:(\mathbb{B}^{2n}(\lambda R),\omega_{0})\hookrightarrow(W,\omega)$ is a symplectic embedding for some $R>0$ and $\lambda>\sqrt{3}$ then $C_{\alpha_{k}}(B_{k}(r))\leq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
To prove the monotonicity and the invariance by conjugation properties one can argue exactly as in Section 4. Also, note that as in Section 4, one deduces immediatly Theorem 1.3 from Theorem 8.1. It thus remains to prove the last two estimates and the proof will mimick the proof we have given in Section 6 with minor modifications.
8.1. Computation of the lower bound
Let us first prove the lower bound that we restate below.
Proposition 8.2.
If $\Psi:(\mathbb{B}^{2n}(R),\omega_{0})\hookrightarrow(W,\omega)$ is a symplectic embedding for $R\in(0,\sqrt{2}]$ then $C_{\alpha_{k}}(B_{k}(r))\geq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
Before proving Proposition 8.2 let us state the following Lemma.
Lemma 8.3.
Let $U$ be an open subset of $(W,\omega)$ then for any $k\in\mathbb{N}_{>0}$
$C_{\alpha_{k}}(\pi_{k}^{-1}(U))\geq C_{\alpha_{1}}(\pi_{1}^{-1}(U)).$
Proof.
For all $\widetilde{\varphi}_{1}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{1}^{-1}(U))$ one can easily construct $\widetilde{\psi}_{k}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{k}^{-1}(U))$ so that the lift of $\widetilde{\psi}_{k}$ to $\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(M_{1},\xi_{1})$ that we denote by $\widetilde{\psi}_{1}$ satisfies $\widetilde{\varphi}_{1}\preceq\widetilde{\psi}_{1}$ and $\widetilde{\psi}_{1}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(\pi_{1}^{-1}(U))$. Moreover thanks to Lemma 7.5 and the monotonicity of $C_{+}^{\alpha_{1}}$ we have
$C_{+}^{\alpha_{k}}(\widetilde{\psi}_{k})\geq C_{+}^{\alpha_{1}}(\widetilde{\psi}_{1})\geq C_{+}^{\alpha_{1}}(\widetilde{\varphi}_{1})$
which concludes the proof. ∎
Proof of Proposition 8.2.
The first 4 steps of the proof will follow the same steps as the proof of Proposition 6.2.
Step 1 : For any $r\in(0,R]$ and $0<\varepsilon<\pi r^{2}$, consider the same function $f:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ as in the first step of the proof of Proposition 6.2 and consider the same map $H:\mathbb{B}^{2n}(r)\to\mathbb{R},\ x\mapsto f(\|x\|^{2})$.
Step 2 : Use the symplectic embedding $\Psi$ to import the construction to $(W,\omega)$, i.e.
$G:W\to\mathbb{R}\quad\quad x\mapsto\left\{\begin{array}[]{ll}H(\Psi^{-1}(x))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(r))\\ 0&\text{ otherwise}\end{array}.\right.$
Step 3 : Import the construction to $(M_{1},\xi_{1})$ by considering the $\alpha_{1}$-Hamiltonian function
$h:M_{1}\to\mathbb{R},\quad\quad\quad x\mapsto G(\pi_{1}(x)).$
Therefore the same arguments as in the proof of Proposition 6.3 shows that for all $T\in[0,1]$
$\mathrm{Spec}^{\alpha_{1}}(\phi_{h}^{T})=\left\{T(\pi r^{2}-\varepsilon)+2\pi m\ |\ m\in\mathbb{Z}\right\}\bigcup\left\{2\pi m\ |\ m\in\mathbb{Z}\right\},$
where $\phi_{h}^{T}\in\mathrm{Cont}_{0}(M_{1},\xi_{1})$.
Step 4 : Therefore arguing as for Proposition 6.3 we get that $c(\widetilde{\varphi}_{T})=T(\pi r^{2}-\varepsilon)$ for all $T\in[0,\delta]$ where $\delta=\min\{1,\frac{2\pi}{\pi r^{2}-\varepsilon}\}$ and $\widetilde{\varphi}_{T}:=[(\phi_{h}^{tT})]$. Since by assumption $R\leq\sqrt{2}$ we deduce that $\delta=1$ and thus $c(\widetilde{\varphi}_{1})=\pi r^{2}-\varepsilon$ which implies that $C_{\alpha_{1}}(B_{1}(r))\geq\pi r^{2}-\varepsilon$. Taking the limit when $\varepsilon$ tends to 00 yields $C_{\alpha_{1}}(B_{1}(r))\geq\pi r^{2}$.
Step 5 : It remains now to apply Lemma 8.3 to deduce that
$C_{\alpha_{k}}(B_{k}(r))\geq C_{\alpha_{1}}(B_{1}(r))\geq\pi r^{2}.$
8.2. Computation of the upper bound
To finish the proof of Theorem 8.1 it thus remains to prove the upper bound that we restate below.
Proposition 8.4.
If $\Psi:(\mathbb{B}^{2n}(\lambda R),\omega_{0})\hookrightarrow(W,\omega)$ is a symplectic embedding for some $R>0$ and $\lambda>\sqrt{3}$ then $C_{\alpha_{k}}(B_{k}(r))\leq\pi r^{2}$ for all $r\in(0,R]$ where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
This proposition will be an adaptation of the proof of Proposition 6.3 together with the following Lemma which is an analogue of Theorem 5.3. First let us remark that
[…]
Lemma 8.5.
Let $U$ be an open subset of $M_{k}$ and $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(M_{k},\xi_{k})$. If $\Pi(\widetilde{\psi})$ $\alpha_{k}$-displaces $U$ then $C_{\alpha_{k}}(U)\leq C_{+}^{\alpha_{k}}(\widetilde{\psi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\psi}).$
Proof.
Let $\widetilde{\varphi}\in\widetilde{\mathrm{Cont}_{0}^{\mathrm{c}}}(U,\xi_{k})$. Then thanks to Proposition 5.5 $C_{+}^{\alpha_{k}}(\widetilde{\psi}\widetilde{\varphi})=C_{+}^{\alpha_{k}}(\widetilde{\psi})$. Moreover since $\widetilde{\psi}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(M_{k},\xi_{k})$ it implies that $\widetilde{\psi}^{-1}\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(M_{k},\xi_{k})$. Hence Corollary 7.3 together with (14) ensure that
$C_{+}^{\alpha_{k}}(\widetilde{\varphi})=C_{+}^{\alpha_{k}}\left(\widetilde{\psi}^{-1}\cdot(\widetilde{\psi}\widetilde{\varphi})\right)\leq C_{+}^{\alpha_{k}}(\widetilde{\psi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\psi}\widetilde{\varphi})=C_{+}^{\alpha_{k}}(\widetilde{\psi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\psi})$
which concludes the proof. ∎
Proof of Proposition 8.4.
Let $r\in(0,R]$, $\delta>0$ be small enough so that $\sqrt{3+2\delta}<\lambda$.
Step 1 : Consider the Hamiltonian function $\widetilde{K}$ supported in $\mathbb{B}^{2n}(\lambda R)$ constructed in the proof of Proposition 6.3
  1. (1)
    $\widetilde{K}$ take values in $[0,r^{2}(\pi+\delta)]$
  2. (2)
    its time $1$ Hamiltonian flow displaces $B_{k}(r)$.
Step 2 : Import the construction to $(W,\omega)$ using $\Psi$, i.e.
$G:W\to\mathbb{R}\quad\quad x\mapsto\left\{\begin{array}[]{ll}H(\Psi^{-1}(x))&\text{ if }x\in\Psi(\mathbb{B}^{2n}(r))\\ 0&\text{ otherwise}\end{array}.\right.$
Step 3 : Import the construction to $(M_{k},\xi_{k})$ using Lemma 6.1. More precisely the $\alpha_{k}$-Hamiltonian function $h:M_{k}\to\mathbb{R}_{\geq 0}$, $x\mapsto G(\pi_{k}(x))$ generates a flow of $\mathbb{T}^{1}$-equivariant contactomorphisms $(\phi_{h}^{t})\subset\mathrm{Cont}_{0}(M_{k},\xi_{k})$. Moreover
  1. (1)
    $h\geq 0$ and $\max h\leq r^{2}(\pi+\delta)$
  2. (2)
    $\phi_{h}^{1}(B_{k}(r))\cap B_{k}(r)=\emptyset$, where $B_{k}(r):=\pi_{k}^{-1}(\Psi(\mathbb{B}^{2n}(r)))$.
Step 4 : Hence thanks to Lemma 8.5
$C_{\alpha_{k}}(B_{k}(r))\leq C_{+}^{\alpha_{k}}(\widetilde{\phi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\phi})$
where $\widetilde{\phi}:=[(\phi_{h}^{t})]\in\widetilde{\mathrm{Cont}_{0}}^{\mathbb{T}^{1}}(M_{k},\xi_{k})$. Finally, since the $\alpha_{k}$-Hamiltonian function $M_{k}\to\mathbb{R}_{\leq 0},\ x\mapsto-h(x)$, generates the contact isotopy $(\phi_{h}^{t})^{-1}$ which represents $\widetilde{\phi}^{-1}$, Proposition 7.2 implies that
$C_{+}^{\alpha_{k}}(\widetilde{\phi}^{-1})+C_{+}^{\alpha_{k}}(\widetilde{\phi})\leq\int_{0}^{1}\max h\mathrm{d}t-\int_{0}^{1}\min h\mathrm{d}t\leq r^{2}(\pi+\delta).$
Letting $\delta$ goes to 00 finishes the proof.